### v1

# Relativistic and Electron Correlation Effects in Static Dipole Polarizabilities for Main-Group Elements

Abstract
In this study, I compute the static dipole polarizability of main-group elements using the finite-field method combined with relativistic coupled-cluster and configuration interaction simulations. The computational results closely align with the values recommended in the 2018 table of static dipole polarizabilities of neutral elements [Mol. Phys. 117, 1200 (2019)]. Additionally, I investigate the influence of relativistic effects and electron correlation on atomic dipole polarizabilities. Specifically, three types of relativistic effects impacting dipole polarizabilities are studied: scalar-relativistic, spin-orbit coupling, and fully relativistic Dirac-Coulomb effects. The results indicate that scalar-relativistic effects are predominant for atoms in Groups 1–2, with minimal influence from spin-orbit coupling effects. Conversely, for elements in Groups 13–18, scalar-relativistic effects are less significant, while spin-orbit coupling significantly affects elements starting from the fourth row in Groups 13–14 and from the fifth row in Groups 15–18. In each category of relativistic effects, the impact of electron correlation is evaluated. The results show that electron correlation significantly influences dipole polarizability calculations, particularly for Groups 1–2 and 13–14 atoms, but is less significant for Groups 15–18 atoms. This study provides a comprehensive and consistent dataset of dipole polarizabilities and contributes to a systematic understanding of the roles of relativistic and electron correlation effects in atomic dipole polarizabilities, serving as a valuable reference for future research.
[…]
University of Stuttgart]Institute of Applied Analysis and Numerical Simulation Numerical Mathematics for High Performance Computing, University of Stuttgart, Pfaffenwaldring 57, 70569, Stuttgart, Germany
1 Introduction
[…]
A recent study by Schwerdtfeger and Nagle compiled an updated table of the most accurate calculated and experimental dipole polarizabilities for neutral atoms in the Periodic Table, with nuclear charges ranging from $Z=1$ to $120$, except for livermorium (Lv, $Z=116$).3 However, there remains potential for improvement in the polarizabilities of open-shell elements and heavy atoms that exhibit significant relativistic effects. This study focuses on all main-group atoms, except for hydrogen (H, $Z=1$), particularly for heavy atoms (i.e., atoms with $Z>40$).
The determination of accurate atomic polarizabilities requires high-level electron correlation treatment within a relativistic framework, particularly for heavy atoms. Over the past two decades, researchers have advanced techniques to incorporate spin-orbit coupling (SOC) effects in dipole polarizability calculations. For closed-shell atoms and open-shell atoms with only one open-shell electron, the coupled cluster theory with single and double substitutions (CCSD) with perturbative triples, CCSD(T), is usually employed within a fully relativistic Dirac-Coulomb framework, where both scalar-relativistic effects and SOC effects can be included. Due to the SOC effect, atoms in Group 14 are also treated as closed-shell atoms. However, for open-shell atoms with more than one open-shell electron in Groups 15–17, the coupled cluster theory is not yet well-developed at the fully relativistic level, and an alternative is multi-reference configuration interaction (MRCI).3
Previous studies have thoroughly investigated the polarizabilities considering SOC effects at the CCSD or CCSD(T) level for cesium (Cs, $Z=55$)4, 5 and francium (Fr, $Z=87$)4, 6, 7 in Group 1. Fleig et al. examined the spin-orbit-resolved static polarizabilities of Group 13 atoms, including gallium (Ga, $Z=31$), indium (In, $Z=49$), and thallium (Tl, $Z=81$), using four-component configuration interaction (CI) and coupled cluster methods with ANO-RCC basis sets in the Dirac-Coulomb framework.8 Borschevsky et al. later improved these results using the Fock space coupled cluster (FSCC) with more extensive Fægri basis sets.4 Pershina et al. first reported the polarizabilities of nihonium (Nh, $Z=113$), computed with the FSCC method and a custom Fægri basis set.9 Dzuba et al. subsequently computed the value for Nh using the SD+CI method in a Dirac framework with Breit interaction and high-order QED correlation, agreeing with the previous work.10 Fleig and Sadlej investigated elements fluorine (F, $Z=9$), bromine (Br, $Z=35$), iodine (I, $Z=53$), and astatine (At, $Z=85$) with two-component spin-orbit coupling CI calculations,11 while tennessine (Ts, $Z=117$) currently only has an empirically estimated value.12 No prior study has provided information on the contribution of the SOC effect to Groups 15–16 atoms, with the exception of moscovium (Mc, $Z=115$), which was examined using the SD+CI approach by Dzuba et al.10 Moreover, errors associated with the recommended values for Groups 15–16 atoms are still significant.3 Currently, no theoretical value exists for Lv.
The main goal of this study is to provide a comprehensive and consistent dataset of dipole polarizabilities for main-group elements. To achieve this, I employed relativistic single-reference coupled cluster (CC) methods and MRCI methods within the Dirac-Coulomb Hamiltonian framework. I included more correlated electrons and virtual orbitals/spinors in correlation energy calculations, using larger basis sets than previous theoretical studies. Additionally, I investigated the impact of electronic correlation and relativistic effects on dipole polarizabilities through a quantitative approach. It should be noted that the data for light atoms ($Z<40$) are also computed in this work. However, their values have been extensively investigated, and improving their theoretical values is beyond the scope of this study. Readers who are interested in these data can refer to Ref. 3 and its latest version.13
The structure of the remainder of this paper is as follows: Section 2 introduces computational methods, followed by computational details in Section 3. Results and discussions are presented in Section 4. Lastly, a summary is given in Section 5. Atomic units are used throughout.
2 Methods
2.1 Relativistic framework
In this study, all energies were calculated within a fully relativistic framework using a four-component relativistic formalism. The relativistic effects can be understood as a combination of SOC effects and scalar-relativistic effects, which account for the contraction or decontraction of the radial electron density.14
2.1.1 Four-component Dirac-Coulomb Hamiltonian
An accurate standard for four-component relativistic electronic-structure theory is based on the Dirac-Coulomb-Breit (DCB) Hamiltonian:15
$\hat{H}_{\text{DCB}}=\sum_{i}\hat{h}_{\text{D}}(i)+\sum_{i<j}\hat{g}_{ij}+\sum_{A<B}V_{AB},$ (1)
where $i$ and $j$ index electrons, $V_{AB}$ is the interaction between nuclei $A$ and $B$, and $\hat{h}_{\text{D}}$ is the one-electron Dirac Hamiltonian. Without external electric fields, it is given by:
$\hat{h}_{\text{D}}(i)=c\bm{\alpha}_{i}\cdot\bm{p}_{i}+c^{2}\beta_{i}+\sum_{A}V_{iA},$ (2)
where $c$ is the speed of light and $\bm{p}$ is the momentum operator. $\bm{\alpha}$ is a vector comprising $\alpha_{x}$, $\alpha_{y}$, and $\alpha_{z}$, which, together with $\beta$, are referred to as the $4\times 4$ Dirac matrices:
$\alpha_{x}=\begin{pmatrix}0_{2}&\sigma_{x}\\ \sigma_{x}&0_{2}\end{pmatrix},\quad\alpha_{y}=\begin{pmatrix}0_{2}&\sigma_{y}\\ \sigma_{y}&0_{2}\end{pmatrix},\quad\alpha_{z}=\begin{pmatrix}0_{2}&\sigma_{z}\\ \sigma_{z}&0_{2}\end{pmatrix},\quad\beta=\begin{pmatrix}I_{2}&0_{2}\\ 0_{2}&-I_{2}\end{pmatrix},$ (3)
where $\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$ are Pauli spin matrices, and $0_{2}$ and $I_{2}$ are $2\times 2$ zero and unit matrices, respectively. The $V_{iA}$ term represents the electron-nuclei interaction on electron $i$ from nucleus $A$.
In the context of two-electron interactions, relativistic corrections up to the second order, specifically the Coulomb-Breit interaction in the Coulomb gauge, are represented by the following equation:16
$\hat{g}_{ij}=\hat{g}^{\text{Coulomb}}+\hat{g}^{\text{Gaunt}}+\hat{g}^{\text{gauge}}=\frac{1}{r_{ij}}-\frac{\bm{\alpha}_{i}\cdot\bm{\alpha}_{j}}{r_{ij}}+\frac{(\bm{\alpha}_{i}\cdot\bm{r}_{ij})(\bm{\alpha}_{j}\cdot\bm{r}_{ij})}{2r_{ij}^{3}},$ (4)
where $\hat{g}^{\text{Coulomb}}$, $\hat{g}^{\text{Gaunt}}$, and $\hat{g}^{\text{gauge}}$ represent the Coulomb interaction, the Gaunt term, and the gauge term, respectively. $\bm{r}_{i}$ ($\bm{r}_{j}$) is the coordinate of electron $i$ ($j$). $\bm{r}_{ij}=\bm{r}_{i}-\bm{r}_{j}$ and $r_{ij}=\|\bm{r}_{ij}\|$. The sum of the Gaunt and gauge terms is also known as the Breit term. When only the Coulomb interaction is considered, the four-component DCH Hamiltonian reduces to the four-component Dirac-Coulomb (DC) Hamiltonian:
$\hat{H}_{\text{DC}}=\sum_{i}\hat{h}_{\text{D}}(i)+\sum_{i<j}\frac{1}{r_{ij}}+\sum_{A<B}V_{AB}.$ (5)
In this study, all calculations are carried out with the DC Hamiltonian.
2.1.2 Exact Two-Component Dirac Hamiltonian
The high computational costs associated with four-component relativistic calculations have prompted the development of methods that use two-component relativistic Hamiltonians. Examples of these methods include the second-order Douglas-Kroll-Hess (DKH) Hamiltonian 17, 18, 19 and zeroth-order regular approximation (ZORA) Hamiltonians.20, 21, 22 However, these methods are all limited to finite orders. Ilias et al. proposed an infinite-order two-component (IOTC) relativistic Hamiltonian, also known as the exact two-component (X2C) method.23 In the X2C method, the relativistic effects can be split into a spin-free part and a spin-dependent part, which correspond to scalar-relativistic and the SOC effects, respectively. In this work, the spin-free X2C method is employed to calculate both non-relativistic and scalar-relativistic properties. More details about the X2C method can be found in recent reviews.16, 24
2.2 Relativistic and Correlation Effects
The Dirac Hartree-Fock (DHF) method was utilized as the uncorrelated method. In this study, the terms “orbital” and “spinor” are used to describe the quantum state of an electron without and with considering SOC effects. Based on these reference states, correlated energies were calculated. Examples of these methods include the second-order Møller-Plesset (MP2) perturbation theory,25 coupled-cluster models CCSD and CCSD(T), where SD stands for single and double excitations, and “(T)” represents perturbed triple excitations,26 and the MRCI method with single and double substitutions (MRCISD).27, 28, 29, 30
In this study, either the four-component (4C) or two-component (2C) Dirac-Coulomb (DC) Hamiltonians are employed to address relativistic effects. The 4C-DC Hamiltonian, which includes Dirac-Coulomb relativistic effects, while the 2C-DC Hamiltonian, implemented in the X2C method, can be used to incorporate more correlated electrons and active virtual orbitals/spinors. The notations “NR-CC”, “SR-CC”, and “DC-CC” are used to refer to CC calculations with non-relativistic, scalar-relativistic, and Dirac-Coulomb relativistic effects, respectively. MRCISD was also used to study the contribution of Dirac-Coulomb relativistic effect, denoted as “DC-CI”, for Group 15–17 elements where the single-reference DC-CC method is not available.
In correlated calculations, the occupied and virtual spaces are usually truncated to reduce computational costs. The orbitals/spinors derived from DHF calculations are then categorized into inner-core, outer-core, valence, and virtual shells. The inner-core electrons remain inactive during the computation of correlation energy, whereas only the electrons in the outer-core and valence shells are active. In practice, outer-core shells with energies lower than the energy cutoff and virtual shells with energies exceeding the cutoff are excluded from the calculation. Thus, one can determine the correlation level of a calculation by considering the number of active electrons and virtual orbitals/spinors. In general, the inclusion of more active electrons and virtual orbitals/spinors leads to more accurate results.
The notation $\alpha_{m}^{n}$ is used to denote dipole polarizabilities computed using a correlated method represented by $m$ and considering a specific relativistic effect denoted by $n$. For instance, $m$ can be one of the following: DHF, MP2, CCSD, CCSD(T), or MRCISD, while $n$ represents NR, SR, or DC, corresponding to non-relativistic, scalar-relativistic, and Dirac-Coulomb relativistic effects, respectively. Given a relativistic effect, the corresponding electron-correlation contribution to polarizabilities is defined as:
$\Delta\alpha_{c}^{n}:=\alpha_{\text{CCSD(T)}}^{n}-\alpha_{\text{DHF}}^{n}$ (6)
where $\Delta\alpha_{c}^{n}$ represents the contribution of electron correlation specific to the relativistic effect denoted by $n$.
Next, the impact of relativistic effects on polarizabilities are estimated. The scalar-relativistic effect on polarizabilities is defined as follows:
$\Delta\alpha_{r}^{\text{SR}}:=\alpha_{\text{CCSD(T)}}^{\text{SR}}-\alpha_{\text{CCSD(T)}}^{\text{NR}}$ (7)
where $\alpha_{\text{CCSD(T)}}^{\text{SR}}$ and $\alpha_{\text{CCSD(T)}}^{\text{NR}}$ are scalar-relativistic and non-relativistic values computed by the CCSD(T) method with the same correlation level.
In addition, the SOC effect on dipole polarizabilities can be then defined as:8, 14
$\Delta\alpha_{r}^{\text{SOC}}=\begin{cases}\alpha_{\text{CCSD(T)}}^{\text{DC}}-\alpha_{\text{CCSD(T)}}^{\text{SR}}&\text{if DC-CC availiable}\\ \alpha_{\text{MRCISD}}^{\text{DC}}-\alpha_{\text{CCSD(T)}}^{\text{SR}}&\text{if DC-CC not availiable}\end{cases}$ (8)
where $\alpha_{\text{CCSD(T)/MRCISD}}^{\text{DC}}$ denotes the DC relativistic polarizabilities computed by the CCSD(T) method, or MRCISD values. While an identical number of active electrons and virtual orbitals/spinors is used for both DC and SR calculations, directly comparing MRCISD values to CCSD(T) data remains challenging due to the contribution from higher-order excitations missed in the MRCISD method. Therefore, Eq. (8) represents both the SOC effects and high excitations when MRCISD values are utilized. To exclusively study the SOC effects, one should compare DC-CI values with MRCISD values that incorporate SR effects, which is beyond the scope of this work. Moreover, the correlation level used in scalar-relativistic calculations might differ between Eqs. (7) and (8). This is because DC calculations at the same correlation level are much more costly than their SR counterparts. Consequently, when the correlation levels differ, notations SRn and SRd are used to refer to the values defined in Eqs. (7) and (8), respectively.
The contribution of full-relativistic DC effects, denoted as $\Delta\alpha_{r}^{\text{DC}}$, can be calculated from the sum of contributions from scalar-relativistic effects and the SOC effect:
$\Delta\alpha_{r}^{\text{DC}}=\Delta\alpha_{r}^{\text{SR}}+\Delta\alpha_{r}^{\text{SOC}}.$ (9)
2.3 Open-shell methods
For the spin-dependent coupling case ($jj$ representation), the averaged polarizability $\bar{\alpha}^{J}$ for the ground state is determined by:14
$\bar{\alpha}^{J}=\frac{1}{2J+1}\sum_{M_{J}}\alpha(J,M_{J})$ (10)
where $\alpha(J,M_{J})$ denotes the polarizabilities for each $J,M_{J}$ component. However, for spin-free cases ($LS$ representation), the equation is analogous to Eq. (10) with $J$ replaced by $L$:14
$\bar{\alpha}^{L}=\frac{1}{2L+1}\sum_{M_{L}}\alpha(L,M_{L})$ (11)
In this work, only $p$-block open-shell elements are studied. The atomic $p$ spinors in the relativistic case are linear combinations of the real $p$ orbitals in the non-relativistic case:11
$\phi\left(\frac{3}{2},\frac{3}{2}\right)$ $=\sqrt{\frac{1}{2}}(p_{x}+ip_{y})|\uparrow\rangle$ $\phi\left(\frac{3}{2},-\frac{3}{2}\right)$ $=\sqrt{\frac{1}{2}}(p_{x}-ip_{y})|\downarrow\rangle$
$\phi\left(\frac{3}{2},\frac{1}{2}\right)$ $=\sqrt{\frac{2}{3}}p_{z}|\uparrow\rangle+\sqrt{\frac{1}{6}}(p_{x}+ip_{y})|\downarrow\rangle$ $\phi\left(\frac{3}{2},-\frac{1}{2}\right)$ $=\sqrt{\frac{2}{3}}p_{z}|\downarrow\rangle+\sqrt{\frac{1}{6}}(p_{x}-ip_{y})|\uparrow\rangle$
$\phi\left(\frac{1}{2},\frac{1}{2}\right)$ $=-\sqrt{\frac{1}{3}}p_{z}|\uparrow\rangle+\sqrt{\frac{1}{3}}(p_{x}+ip_{y})|\downarrow\rangle$ $\phi\left(\frac{1}{2},-\frac{1}{2}\right)$ $=\sqrt{\frac{1}{3}}p_{z}|\downarrow\rangle-\sqrt{\frac{1}{3}}(p_{x}-ip_{y})|\uparrow\rangle,$ (12)
[…]
2.3.1 Group 13 elements
The spin-orbit components of Group 13 atomic states can be represented as the $p^{1}$ states using atomic spinors $\phi(j,m_{j})$ due to the presence of only one electron in $p$-spinor. In the ground state, only the spinors $\phi\left(\frac{1}{2},\pm\frac{1}{2}\right)$ are occupied. When an external perturbation, such as an electric field along the $z$ axis, is applied, the two degenerate spinors $\phi\left(\frac{3}{2},\pm\frac{3}{2}\right)$ and spinors $\phi\left(\frac{3}{2},\pm\frac{1}{2}\right)$ split, resulting in a non-degenerate state. The energy shift is dependent on the $p_{z}$-component and generally, spinors with a larger $p_{z}$ contribution will experience a larger energy shift than those with a smaller contribution.11 However, as weak electric fields are typically applied, the energy shift is much smaller in absolute value than the spin-orbit splitting. As a result, the $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ states will have a smaller dipole polarizability than the $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ states.
Based on the analysis in Ref. 8, the value $\bar{\alpha}_{J=3/2}-\alpha_{J=1/2}$ can be interpreted as the effect of both SOC and the contraction (decontraction) of the $p_{1/2}$ ($p_{3/2}$) spinor in total. The SOC effect can also be determined by comparing the $M_{L}$-averaged polarizability, calculated using SR-CC methods, to the value of the $J=\frac{1}{2}$ ground state computed using the DC-CC method. Moreover, the DC-CI method allows for the computation of excited-state polarizabilities, facilitating the study of spin-orbit resolved polarizabilities for these excited state components.
2.3.2 Group 14 elements
As previously discussed, the $J=\frac{1}{2},M_{J}=\frac{1}{2}$ spinors have a lower energy than the $\phi\left(\frac{3}{2},m_{j}\right)$ spinors, resulting in the full occupation of the $\phi\left(\frac{1}{2},m_{j}\right)$ spinors. This means that Group 14 atoms can be considered as closed-shell atoms due to the SOC effect.
2.3.3 Group 15 elements
The DC-CI calculations were performed to investigate the electron-correlation and the SOC effect on polarizabilities. Additionally, the polarizabilities of different $M_{J}$ sub-states were studied. In practice, DC-CI eigenvectors are often represented by string occupation and the spin-projection value $M_{K}$, which is an integer multiple of $\frac{1}{2}$ for a given determinant.11 As a result, the different components of ${{}^{4}}P_{3/2}$ are not pure states, and may be combined with high $M_{K}$ states that have the same $M_{J}$. To address this, the method proposed by Fleig et al. 11 was used to approximately specify the $M_{J}$ state. However, since the primitive determinant is determined by $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$ for $M_{K}=\frac{1}{2}$, the polarizabilities of these states are thus approximately considered to represent ground-state values. In this work, the states $J=\frac{3}{2},M_{J}=\pm\frac{1}{2}$ refer to the states with lower energy compared to the states $J=\frac{3}{2},M_{J}=\pm\frac{3}{2}$ when an external electric field is applied along the $z$-axis.
2.3.4 Group 16 elements
The polarizabilities of different $M_{J}$ components of the ${{}^{3}}P_{2}$ states were determined using a similar procedure to that used for Group 15 elements. The DC-CI method was used to calculate the polarizabilities for states $M_{J}=0$, $M_{J}=\pm 1$ and $M_{J}=\pm 2$, and the $M_{J}$-averaged polarizability is approximated as the ground-state value. In this work, the states $J=2,M_{J}=\pm 1$ refer to the states with lower energy compared to the states $J=2,M_{J}=\pm 2$ when an external electric field is applied along the $z$-axis.
2.3.5 Group 17 elements
The spin-orbit components of the ground states of halogen atoms can be represented as the inverse of the $p^{5}$ hole states by using atomic spinors $\phi(j,m_{j})$ as there is only one hole in the $p$-spinor for these atoms.11 The polarizabilities of different components of the ${{}^{2}}P_{3/2}$ states were determined using a similar procedure to that used for Group 15 elements. The DC-CI method was used to calculate the polarizabilities for states $M_{J}=\pm 3/2$ and $M_{J}=\pm 1/2$, and the $M_{J}$-averaged polarizability is approximated as the ground-state value. Moreover, the computational values for states ${{}^{2}}P_{1/2}$ are also computed.
2.4 Finite-field methods
The static dipole polarizabilities presented in this work were computed using finite-field methods. 31. The energy of an atom in a homogeneous electric field of strength $F_{z}$ along the $z$-axis is represented as follows:31
$E(F_{z})=E_{0}-\frac{1}{2}\alpha F_{z}^{2}-\frac{1}{4!}\gamma F_{z}^{4}-\frac{1}{6!}\gamma_{2}F_{z}^{6}-\cdots$ (13)
where $E_{0}$ denotes the field-free energy, while $\alpha$, $\gamma$, and $\gamma_{2}$ correspond to the dipole polarizability, the dipole hyperpolarizability, and the second dipole hyperpolarizability, respectively. For the purpose of this study, the expansion is truncated at the third term, allowing one to calculate the $\alpha$ and $\gamma$:
$E(F_{z})\approx E_{0}-\frac{1}{2}\alpha F_{z}^{2}-\frac{1}{4!}\gamma F_{z}^{4}.$ (14)
The least-squares procedure is used instead of numerical differentiation for reasons discussed in Ref. 32. In practice, this procedure could lead to an unreasonable $\gamma$, including extremely positive or negative values. In these cases, an expansion only including $\alpha$ is also used:
$E(F_{z})\approx E_{0}-\frac{1}{2}\alpha F_{z}^{2}.$ (15)
The results obtained from Eq. (14), denoted as $\alpha(2)$, should always be used instead of the values obtained from Eq. (15), denoted as $\alpha(1)$, except in cases where all the following conditions are satisfied:
  1. 1.
    The relative difference between $\alpha(1)$ and $\alpha(2)$, denoted as $\tau$, is greater than 0.6%, i.e.,
    $\tau=\frac{|\alpha(2)-\alpha(1)|}{\alpha(2)}\times 100\%>0.6\%.$ (16)
  2. 2.
    An invalid $\gamma$ is obtained from Eq. (14). For simplicity, in this work, an invalid $\gamma$ is defined as when at least one of the following conditions is satisfied:
    1. (a)
      For CCSD(T) methods,
      $\gamma_{\text{CCSD(T)}}$ $<0,$ $\gamma_{\text{CCSD}}$ $<0,$ $\epsilon=\frac{\max(\gamma_{\text{CCSD(T)}},\gamma_{\text{CCSD}})}{\min(\gamma_{\text{CCSD(T)}},\gamma_{\text{CCSD}})}$ $>10,$ (17)
      where $\gamma_{\text{CCSD(T)}}$ and $\gamma_{\text{CCSD}}$ are the $\gamma$ values obtained by the CCSD(T) and CCSD methods, respectively. Here, $\max(a,b)$ represents the maximum value between $a$ and $b$, and $\min(a,b)$ represents the minimum value between $a$ and $b$. $\epsilon$ represents the ratio of the maximum to the minimum of these $\gamma$ values, indicating a significant difference between the $\gamma$ values from the two methods when $\epsilon>10$.
    2. (b)
      For MRCISD methods:
      $\gamma_{\text{MRCISD}}$ $<0,$ (18)
    where $\gamma_{\text{MRCISD}}$ refers to the $\gamma$ value obtained by the MRCISD method.
2.5 Uncertainty estimation
The uncertainty estimation is analyzed empirically based on the composite scheme, which has been widely employed in previous studies to obtain the final value of an atomic property, denoted as $P_{\text{final}}$.33, 14, 34 It should be noted that the $P_{\text{final}}$ values are not provided in this work due to their empirical nature and are only used for uncertainty estimation empirically.
2.5.1 Composite scheme of the CC method
Generally, in the CC calculations,
$P_{\text{final}}=P_{\text{CCSD}}+\Delta P_{\text{basis}}+\Delta P_{\text{(T)}}\text{ or }\Delta P_{\text{T}}+\Delta P_{\text{Q}}+\Delta P_{\text{core}}+\Delta P_{\text{vir}}+\Delta P_{\text{fitting}}+\Delta P_{\text{others}}$ (19)
with
$\Delta P_{\text{(T)}}$ $=P_{\text{CCSD(T)}}-P_{\text{CCSD}}$ (20)
$\Delta P_{\text{T}}$ $=P_{\text{CCSDT}}-P_{\text{CCSD}}$ (21)
$\Delta P_{\text{Q}}$ $=P_{\text{CCSDTQ}}-P_{\text{CCSDT}}$ (22)
where $P_{\text{CCSD}}$, $P_{\text{CCSD(T)}}$, $P_{\text{CCSDT}}$, and $P_{\text{CCSDTQ}}$ are the values calculated for the property using CCSD, CCSD(T), CCSDT (CC singles, doubles, and triples), and CCSDTQ (CC singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$.35, 34 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual orbitals/spinors.35 $P_{\text{CCSD}}$ or $P_{\text{CCSD(T)}}$ is the value for property $P$ using the most accurate CCSD or CCSD(T) method in this work. $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). For completeness, $\Delta P_{\text{others}}$ corresponds to all other corrections, e.g., the difference between $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{T}}$ when $\Delta P_{\text{(T)}}$ is used, or other relativistic effects, including Gaunt, Breit, or QED interactions.
It should be noted that the most accurate calculation in CC used in this work is CCSD(T). When CCSDT and CCSDTQ values are available in the literature, $\Delta P_{\text{T}}$ and $\Delta P_{\text{Q}}$ are employed; otherwise, $\Delta P_{\text{(T)}}$ is used, and $\Delta P_{\text{Q}}$ is neglected with the inclusion of a larger empirical error. Also, $\Delta P_{\text{others}}$ is included when its valid value is available in the literature.
2.5.2 Composite scheme of the MRCI method
In the CI calculation,
$P_{\text{final}}=P_{\text{MRCISD}}+\Delta P_{\text{basis}}+\Delta P_{\text{T}}+\Delta P_{\text{Q}}+\Delta P_{\text{core}}+\Delta P_{\text{vir}}+\Delta P_{\text{fitting}}+\Delta P_{\text{others}}$ (23)
with
$\Delta P_{\text{T}}$ $=P_{\text{MRCISDT}}-P_{\text{MRCISD}}$ (24)
$\Delta P_{\text{Q}}$ $=P_{\text{MRCISDTQ}}-P_{\text{MRCISDT}}$ (25)
where $P_{\text{MRCISD}}$, $P_{\text{MRCISDT}}$, and $P_{\text{MRCISDTQ}}$ are the values calculated for the property using MRCISD, MRCISDT (MRCI singles, doubles, and triples), and MRCISDTQ (MRCI singles, doubles, triples, and quadruples), respectively. $\Delta P_{\text{basis}}$ is the error due to the use of the finite-size basis functions for computing $P_{\text{MRCISD}}$.35, 34 $\Delta P_{\text{core}}$ is the correction due to the more outer-core shells, while $\Delta P_{\text{vir}}$ is the possible error due to the neglected contributions from the frozen high-lying virtual spinors.35 $\Delta P_{\text{fitting}}$ is the numerical error resulting from the fitting procedure using Eq. (14) or Eq. (15). Similarly, $\Delta P_{\text{others}}$ corresponds to all other corrections as in CC calculations.
It should be noted that the highest-order calculation in CI is MRCISD in this work. When MRCISDT and MRCISDTQ values are available in the literature, $\Delta P_{\text{T}}$ and $\Delta P_{\text{Q}}$ are employed; otherwise, they are neglected with a larger empirical error used.
2.5.3 Errors estimation
Based on the composite schemes, the possible error sources causing the uncertainty in $P_{\text{final}}$ in this work are similar to those proposed in Refs. 14, 35, 34:
  1. 1.
    The first one is due to the finite basis sets used for computing $P_{\text{CCSD}}$/$P_{\text{CCSD(T)}}$ and $P_{\text{MRCISD}}$, i.e., $\Delta P_{\text{basis}}$. Since the errors converge quickly, the errors in these values are taken as half of the differences of $P$ values obtained by “d-aug-dyall.v3z” and “d-aug-dyall.v4z”, corresponding to the two most diffuse exponents for the respective atomic shells in the original basis sets “dyall.v3z” and “dyall.v4z”, respectively.14 In this work, all most accurate calculations of $p$-block elements are carried out using “dyall.acv4z”, where one more diffuse but more core basis functions are added to the original “dyall.v4z” basis set. See Ref. 36 for more details. This encourages the belief that the results in this work are also closer to the complete basis sets. For simplicity, the same empirical estimate method defined in Ref. 14 is used when results with additional basis functions are available in this work. Otherwise, a conservative choice is to include the full difference between this work and the most accurate results obtained from the literature as the error.
  2. 2.
    The second error is from $\Delta P_{\text{(T)}}$, $\Delta P_{\text{T}}$, or $\Delta P_{\text{Q}}$. Half of them is also taken into account as the error as in Ref. 14.
  3. 3.
    The third one is from $\Delta P_{\text{core}}$ due to the number of outer-core shells included in the calculations.
  4. 4.
    The fourth one is from $\Delta P_{\text{vir}}$ due to neglecting contributions from the outer-core shells.
  5. 5.
    The next one is from $\Delta_{\text{fitting}}$ due to the numerical fitting procedure. Specifically, the standard errors of the regression coefficients are evaluated using the residuals from the least-square solutions, 37 which has been implemented in Ref. 38.
  6. 6.
    The last one is the error due to $\Delta P_{\text{others}}$. It is usually computationally demanding to compute them systematically in practice. For simplicity, it is only included when available in the previous literature.
The uncertainty of the final $P_{\text{final}}$ values is determined by taking the quadrature of the $\Delta P_{\text{basis}}$, $\Delta P_{\text{(T)}}$, $\Delta P_{\text{Q}}$, $\Delta P_{\text{core}}$, $\Delta P_{\text{vir}}$, $\Delta P_{\text{fitting}}$, and $\Delta P_{\text{others}}$ contributions. More details on the uncertainty estimation for each element will be presented in Section 4.
3 Computational Details
In this section, the computational details are introduced, including the computational basis sets used and the selected correlation level.
3.1 Basis sets
In this work, I mainly used the Dyall quadruple-$\zeta$ family uncontracted basis sets.39, 40, 41, 42, 43, 44, 45, 46 For Group 1–2 atoms, I used ANO-RCC basis set 47, 48 instead of the Dyall basis set. The original ANO-RCC basis sets are augmented with additional functions, extending each type of function in an even-tempered manner. The exponential coefficients of the additional augmented functions were determined using the equation $\zeta_{N+1}=\zeta_{N}^{2}/\zeta_{N-1}$, where $\zeta_{N}$ and $\zeta_{N-1}$ represent the smallest exponents for each atomic shell in the default basis sets.14 These new basis sets are labeled as “s-aug-ANO-RCC” and “d-aug-ANO-RCC” for single and double augmentations, respectively. After test calculations, the “s-aug-ANO-RCC” basis set was selected for Group 1 elements. For Group 2 elements, the “s-aug-dyall.cv4z” basis set was used for Sr, Ba, and Ra, while the “d-aug-dyall.cv4z” basis set was utilized for Be, Mg, and Ca. In addition, “dyall.cv4z” and “d-aug-dyall.cv4z” are also used in test calculations to assess the error due to finite basis sets, which will be discussed later. For $p$-block elements, I used the “dyall.acv4z” basis set, which includes one additional diffuse and more core basis functions compared to the original “dyall.v4z” basis set.
3.2 Correlation level
All calculations are performed using the DIRAC18 package.49 In general, orbitals within an energy range of -20 to 25 a.u. are correlated in CC and CI calculations. The convergence criterion for the CC method was set at $10^{-10}$. In SR-CC calculations, any deviation between SRn and SRd values was observed to be less than 1% in this work. For the MRCISD calculations, the energy criterion for convergence was set to $10^{-13}$. Further computational details for $p$-block open-shell elements are provided in the following subsections.
3.2.1 Group 13 elements
DHF orbitals, used for DC-CC calculations, are optimized by averaging over various $np$ occupations with one open-shell electron. The open shells are restricted to the $np_{1/2}$ spinors, leading to a single determinant reference as detailed in previous work 8. Additionally, DC-CI calculations were employed to obtain excited-state components, with DHF orbitals constructed from averaging over all $np_{1/2}$ and $np_{3/2}$ spinors.
3.2.2 Group 14 elements
In DC calculations, Group 14 elements are typically considered as closed-shell systems. Consequently, the 2C-DC Hamiltonian was used in all DC-CC calculations for Group 14 elements, allowing for the inclusion of more correlated electrons. However, this approach is not applicable to the carbon atom, as the $\phi\left(\frac{1}{2},m_{j}\right)$ spinor occupied state does not constitute a suitable reference wave function, which can lead to convergence issues. Therefore, the DC calculation for the carbon atom was conducted using the 4C-DC-CC method.
3.2.3 Group 15–17 elements
Only DC-CI calculations were performed to explore the effects of electron correlation and the contribution of the SOC effect on polarizabilities. The multiple-reference wavefunction is used in the DHF method of DC-CI calculation, while the single-reference wavefunction is used in DHF of SR-CC calculation. Therefore, DHF energies obtained in DC-CI calculations are not employed here to compute properties related to $\alpha_{\text{DHF}}$. As a result, $\Delta\alpha_{c}^{\text{SOC}}$ is not available for Group 15–17 elements.
3.3 Finite-field method
Electric fields with strengths of 0.000, 0.0005, 0.001, 0.002, and 0.005 a.u. are applied to each element in the calculation of dipole polarizabilities, with the exception of Group 1 elements. For these elements, a new set of uniform electric fields with strengths of 0.000, 0.0005, 0.001, 0.0015, and 0.002 a.u. is applied due to their large polarizabilities. For scalar-relativistic CC methods, a larger range of finite electric fields, ranging from 0.000 to 0.020 a.u. in increments of 0.001, is incorporated to obtain more accurate results for $p$-block elements. The specified number of electric fields for each atom can be found in the Supporting Information. The energies are then fitted to Eq. (13) using a least-squares method to obtain the dipole polarizability.
4 Results
In this section, I present the computational values for main-group elements and discuss the impact of relativistic effects and electron correlation on dipole polarizabilities. All results related to $\alpha$ obtained by fitting Eq. (14) are provided in Tables S1-S8 in the Supporting Information. Results for $\alpha$ obtained by fitting Eq. (15) can be found in Tables S9-S16 in the Supporting Information. It was observed that for the most accurate DC values, only aluminum (Al, $Z=13$) should be updated using Eq. (15) when $\tau>0.6\%$. Additionally, Al and oganesson (Og, $Z=118$) should be updated using Eq. (15) when $\tau>0.5\%$. For $\tau>0.1\%$, the DC values of potassium (K, $Z=19$), rubidium (Rb, $Z=37$), Al, Nh, and Og should be computed using Eq. (15). Therefore, unless otherwise specified, the results derived from Eq. (14) are consistently utilized in the following discussions.
4.1 Group 1 elements
4.1.1 Comparison with literature
Table 1 summarizes the CCSD(T) results for Group 1 elements, compared with the recommended values from Ref.3. Additional results can be found in Table S1. The results obtained by Eq. (15) can be found in Table S9, and the corresponding summary of the most accurate results is listed in Table S17.
Table 1: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 1 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
Li Na K Rb Cs Fr
$\hat{H}$ State Method
NR ${}^{2}S$ CCSD(T) $164.34$ $163.98$ $295.56$ $344.65\pm 0.01$ $474.79\pm 0.05$ $541.62\pm 0.04$
SR ${}^{2}S$ CCSD(T) $164.29$ $162.99$ $289.96$ $319.34\pm 0.01$ $397.40\pm 0.02$ $319.36\pm 0.01$
DC ${}^{2}S_{1/2}$ CCSD(T) $164.31\pm 0.01$ $163.27\pm 0.09$ $290.64\pm 0.24$ $320.08\pm 0.29$ $398.24\pm 0.41$ $317.40\pm 0.57$
Rec. --- - --- - $164.31\pm 0.20$ $163.27\pm 0.81$ $289.40\pm 2.13$ $320.08\pm 0.83$ $398.24\pm 2.71$ $317.40\pm 0.99$
Ref. 3 --- - --- - $164.11$ $162.7\pm 0.5$ $289.7\pm 0.3$ $319.8\pm 0.3$ $400.9\pm 0.7$ $317.8\pm 2.4$
In addition, the most accurate theoretical and experimental values for the entire periodic table have been summarized in Refs. 3 and 13. These values for Group 1 elements are collected in Table S25 to validate the results in this work. Several observations can be made and discussed as follows. In this work, only computational and experimental values close to the recommended values in Ref. 3 are compared.
The DC result ($164.31\pm 0.01$) for lithium (Li, $Z=3$) aligns closely with experimental results ($164.0\pm 3.4$ 50 and $164.2\pm 1.1$ 51). The most accurate computational value ($164.1125\pm 0.0005$ 52) is taken as the recommended value in Ref. 3. The difference between the recommended value and the result in this work is only 0.12%, which could be attributed to the absence of other relativistic effects, such as Breit or QED. It should be noted that the full relativistic contribution to $\alpha$ is $-0.07$ a.u. at both the DHF and CCSD levels, as shown in Table S1, which agrees with the value ($-0.07$ a.u.) obtained by Lim et al. at the DHF level.53
For sodium (Na, $Z=11$), the deviation between the DC result ($163.27\pm 0.09$) and the recommended value ($162.7\pm 0.5$) in Ref.3, based on the experimental result by Ekstrom et al., is less than 0.36%.54 Notably, the lower bound of the DC result is slightly larger than the upper bound of both the recommended value and another experimental value ($162.7\pm 0.1$) reported by Holmgren et al.55 However, the DC result for Na ($163.27\pm 0.09$) is close to the experimental value ($161\pm 7.5$) reported by Ma et al.56 Additionally, the DC result is slightly larger than the recommended computational value (162.88) in Refs. 57, 58, obtained using a composite scheme by Thakkar and Lupinetti. The NR and SR values of Na (163.98 and 162.99, respectively) in this work are in good agreement with the computational values of 163.90 and 162.90, respectively, as reported by Thakkar and Lupinetti.57, 58 Therefore, the primary difference between the DC value and the final recommended value in Refs. 57, 58 lies in the method of incorporating the SOC contribution to $\alpha$. Specifically, in Refs. 57, 58, a value of $-0.02$ a.u. evaluated at the DHF level from Ref. 53 is used as the SOC contribution and is added to the SR value. In contrast, this work evaluates the SOC contribution as $+0.28$ a.u. at the CCSD(T) level. It should be noted that in this study, the SOC contributions at the DHF and CCSD levels are $-0.04$ and $-0.03$ a.u., respectively. Furthermore, this study shows that the NR value also agrees well with the recent non-relativistic CCSD(T) value obtained by Smiałkowski and Tomza (163.9 a.u.) at CCSD(T)/aug-cc-pwCV5Z,59 suggesting the reliability of the method used in this work.
For K, the lower bound of the DC value ($290.64\pm 0.24$) is slightly higher than the upper bound of the recommended value ($289.7\pm 0.3$),3 which is based on experimental values.60, 61 However, the DC CCSD(T) value obtained using Eq. (15) is $289.40\pm 0.20$, as shown in Table S17, which agrees well with the recommended value in Ref. 3. Therefore, one reason for the discrepancy could be that a negative $\gamma$ obtained from Eq. (14) is invalid, which is observed when $\tau>0.1\%$ is used in Eq. (16).
Moreover, the DC value ($290.64\pm 0.24$) is in good agreement with other experimental values ($292.9\pm 6.1$ 50 and $290.6\pm 1.4$ 55). Compared to the computational values, the DC result is slightly higher than the RLCCSD value (289.1),62 where the linear approximation CCSD method is used. Additionally, the DC value ($290.64\pm 0.24$) agrees well with the recent result ($290.30\pm 0.23$) computed by the relativistic all-order method.63
The SR result (289.96) is in good agreement with the DK-CCSD(T) result ($291.1\pm 1.5$) reported by Lim et al. using all-electron calculations.64 Furthermore, the SR value (289.96) is also close to the recent result (289.6) computed by CCSD(T) with relativistic energy-consistent pseudopotentials (ECP).59
The most accurate DC value for Rb in this work is $320.08\pm 0.29$ a.u., which is in good agreement with previous experiments ($319\pm 6$ 50, $318.8\pm 1.4$ 55, $319.8\pm 0.3$ 60, 61, $319.2\pm 6.1$ 57) and the recommended value ($319.8\pm 0.3$).3 The DC value is generally higher than recent values (318.5±0.6plus-or-minus318.50.6318.5$\pm$ 0.6318.5 ± 0.665, 318.38±0.38plus-or-minus318.380.38318.38$\pm$ 0.38318.38 ± 0.3863) obtained by the relativistic all-order method and its variants, but it is in good agreement with the latest result ($319.5\pm 1.5$)66, which includes other relativistic effects, e.g., Breit and QED, etc. It should be noted that the DC value obtained using Eq. (15) is $319.08\pm 0.17$, which is less than the recommended value in Ref. 3. This suggests that a more accurate $\gamma$ should be used to obtain $\alpha$ for Rb. Moreover, the SR value ($319.34\pm 0.01$) is slightly larger than the latest result (317.4) computed by scalar-relativistic CCSD(T) with the ECP method.59
The DC value for Cs ($398.24\pm 0.34$) is slightly lower than the value from previous theoretical work, which was 399.0, reported by Borschevsky et al.,4 computed using the CCSD(T) method with the Fægri basis set.67 The discrepancy could be attributed to the different basis set and fitting procedure used in this study. To study the impact of basis sets, I used the same Fægri set as in Ref. 4. The virtual orbitals with energies above 40 a.u. were omitted instead of 50 a.u. in the previous work to reduce computational cost. It should be noted that five data points were used to fit $\alpha$ defined in Eq. (14) using the least-squares procedure, instead of the three data points used with numerical differentiation in Ref. 4. The result ($399.52\pm 0.62$ a.u.), as shown in Table S1, was closer to 399.0 in Ref. 4. It is also in excellent agreement with the recent results (399.9±0.6plus-or-minus399.90.6399.9$\pm$ 0.6399.9 ± 0.665 and 399.74±0.55plus-or-minus399.740.55399.74$\pm$ 0.55399.74 ± 0.5563) computed by relativistic all-order methods. When Eq. (15) is used, the results from the least-squares procedure are $400.51\pm 0.2$ and $400.60\pm 0.37$ a.u. for the Fægri and s-aug-ANO-RCC basis sets, respectively. These results are closer to the recommended value ($400.9\pm 0.7$ a.u.) from Ref. 3 compared to those obtained by fitting Eq. (14). The larger difference in DC CCSD(T) values obtained with s-aug-ANO-RCC between using Eqs. (14) and (15) can be attributed to a larger $\gamma$ obtained compared to the Fægri basis. Therefore, simply removing $\gamma$ and higher-order contributions in Eq. (15) could lead to a higher numerical error for Cs. Surprisingly, this error is even as significant as the SOC contribution to $\alpha$, as shown in Table 1. The SR value ($397.40\pm 0.02$) in this work is much higher than the result (391) obtained by scalar-relativistic CCSD(T) with the ECP method.59 One reason could be the pseudopotential used to account for the scalar-relativistic effect in Ref. 59.
The DC value of Fr ($317.40\pm 0.57$) is in excellent agreement with the recommended value ($317.8\pm 2.4$) reported in Ref. 3 and the recent computational values ($317.1\pm 1.3$,7 $316.6\pm 2.4$,68 316.6±1.5plus-or-minus316.61.5316.6$\pm$ 1.5316.6 ± 1.563) using the relativistic all-order method and its variants. There is a significant difference between the results of the current work and those reported in Ref. 4, where the Fægri basis set and fewer data points were used. To investigate this discrepancy, I chose the same data points as in Ref. 4 and fitted Eq. 15. The resulting value is 318.34, which is still much larger than the value reported in Ref. 4. Therefore, the primary reason for this discrepancy could be the different basis functions used in the two studies.
4.1.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 1 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are not estimated explicitly. They are included implicitly in the error due to $\Delta P_{\text{basis}}$. Several SR test calculations were performed for basis sets using ANO-RCC and d-aug-ANO-RCC, compared to the results with s-aug-ANO-RCC.
In SR-CC calculations of Li, as shown in Table S1, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 164.30, 164.29, and 164.28 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.01 a.u., with half of it taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 0.07 a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$. To be conservative, I take the difference between this study and the recommended value as the error, i.e., 0.20 a.u., which is mainly due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.21 a.u. in this study for Li is acceptable.
In SR-CC calculations of Na, as shown in Table S1, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 163.00, 162.99, and 162.97 a.u., respectively, suggesting that the error due to using s-aug-ANO-RCC basis set is about 0.02 a.u., with half of it taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 1.11 a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$. To be conservative, the difference between this study and the recommended value is taken as the error, i.e., 0.57 a.u., which is mainly due to neglecting other relativistic and higher-order correlation effects, i.e., $\Delta P_{\text{others}}$. In conclusion, the error of 0.80 a.u. in this study for Na is acceptable.
In SR-CC calculations of K, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 284.74, 289.96, and 290.93 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.97. To be conservative, it is taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 12.19 a.u., which is much higher than the difference between the recommended value and the DC result in this work, i.e., 0.94 a.u. To be conservative, I take twice the latter difference as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 2.12 a.u. in this study for K is acceptable.
In SR-CC calculations of Rb, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 314.32, 319.34, and 319.89 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.55 a.u., which is therefore taken as the error due to finite basis sets. To be conservative, I take twice the difference between the recommended value and the DC results in this work, 0.56 a.u. in total, as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.78 a.u. in this study for Rb is acceptable.
For Cs, I performed DC-CC calculations with d-aug-ANO-RCC, and the CCSD(T) result is $398.16\pm 0.33$, which agrees well with the value evaluated at the CCSD(T)/s-aug-ANO-RCC level ($398.24\pm 0.34$). The difference of 0.08 between these two basis sets suggests that the result is converged with the ANO-RCC basis set for $\alpha$ calculations. I therefore take the difference between the results with s-aug-ANO-RCC and Fægri basis sets as the error due to the finite basis sets ($\Delta P_{\text{basis}}$), i.e., 1.28 a.u. Based on the calculations with different fitting procedures used, the numerical errors of 2.35 a.u. are added as the error due to other effects. In conclusion, the error of 2.68 a.u. in this study for Cs is acceptable.
In SR-CC calculations of Fr, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are $318.72$, $319.36$, and $319.21$ a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.15 a.u., which is taken as the error due to finite basis sets. To be conservative, I take twice the difference between the recommended value and the DC results in this work (0.40 a.u.) as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.81 a.u. in this study for Fr is acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 1. It should be noted that for K, the DC CCSD(T) result obtained using Eq. (15) is taken as the recommended value, as shown in Table S17.
4.1.3 The correlation and relativistic effects on polarizabilities
Figure 1(a) demonstrates that non-relativistic polarizabilities increase monotonically with atomic numbers, starting from Na. In contrast, the relativistic data increase with increasing atomic numbers from Na to Cs, but decrease from Cs to Fr, with the value for Fr being close to that of Rb. This trend confirms the irregular progression of values with atomic numbers for Group 1 elements, driven by the relativistic contraction of the valence $s$ shell.53, 64 The polarizability of Na is slightly lower than that of Li due to reduced screening of the nuclear charge by the $p$ electrons in Na, resulting in more compact valence shells and lower polarizabilities.
Figure 1(b) indicates that the scalar-relativistic effect is more pronounced than the SOC effect, leading to a roughly $Z^{2}$ increase in relativistic effects on dipole polarizabilities due to the scalar-relativistic contribution. The SOC contribution to dipole polarizabilities for Group 1 atoms is small; generally, the SOC effect slightly increases the polarizabilities of Group 1 elements, except for Fr, where the SOC effect reduces dipole polarizability by nearly 2. This observation aligns with previous findings.53, 64
As depicted in Figure 1(c), at the non-relativistic level, the electron correlation contribution increases monotonically in absolute value with increasing atomic numbers. In contrast, relativistic calculations exhibit the same trend up to Cs, although the contributions from Rb and Cs are slightly lower than their non-relativistic counterparts. An unusual trend is observed for Fr, where the electron correlation contributions in relativistic calculations are significantly lower than those in the non-relativistic case.
Figure 1: Polarizabilities (in a.u.) of Group 1 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.2 Group 2 elements
4.2.1 Comparison with literature
Table 2 summarizes the CCSD(T) results for Group 2 elements, compared with the recommended values from Ref. 3. Additional results can be found in Table S2. The results obtained by Eq. (15) can be found in Table S10, and the corresponding summary of the most accurate results is listed in Table S18.
Table 2: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 2 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
Be Mg Ca Sr Ba Ra
$\hat{H}$ State Method
NR ${}^{1}S$ CCSD(T) $37.80$ $71.54$ $159.90$ $209.64\pm 0.01$ $316.58\pm 0.01$ $378.82\pm 0.01$
SR ${}^{1}S$ CCSD(T) $37.79$ $71.20$ $157.61$ $197.82$ $273.96$ $247.69\pm 0.01$
DC ${}^{1}S_{0}$ CCSD(T) $37.78$ $71.19$ $157.55$ $197.67\pm 0.01$ $273.51$ $245.68\pm 0.01$
Rec. --- - --- - $37.78\pm 0.10$ $71.19\pm 0.32$ $157.55\pm 1.42$ $197.67\pm 2.31$ $273.51\pm 4.15$ $245.68\pm 4.91$
Ref. 3 --- - --- - $37.74\pm 0.03$ $71.2\pm 0.4$ $160.8\pm 4$ $197.2\pm 0.2$ $272.9\pm 10$ $246.0\pm 4$
In addition, the most accurate theoretical and experimental values for the entire periodic table have been summarized in Refs. 3 and 13. These values for Group 2 elements are collected in Table S26 to validate the results in this work.
The DC values are in good agreement with the recommended values, with the exception of beryllium (Be, $Z=4$) and strontium (Sr, $Z=38$). For these elements, the DC values slightly exceed the upper bounds of the recommended values. Next, a detailed comparison between this work and the results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared.
The DC result (37.78) with CCSD(T) for Be aligns closely with recent computational theoretical values ($37.73\pm 0.04$ 69, 70 and 37.787 71) computed by the relativistic configuration interaction plus core polarization (RCICP) method, as well as the result ($37.76\pm 0.22$) obtained by CI+MBPT2 with experimental data.72 The DC CCSD value (37.94), as shown in Table S2, is also in good agreement with the results ($37.80\pm 0.47$ 73 and $37.86\pm 0.17$ 74) obtained by relativistic CCSD. The most accurate SR and NR values (37.79 and 37.80) in this work are both higher than the corresponding results (37.69 and 37.71) in Ref. 75, where CCSD(T) calculations were performed at the complete basis set limit, suggesting that the difference could be mainly attributed to the finite basis sets used in this work. Additionally, the NR value is also close to other computational values (37.755,76 37.75,75 37.7,59 and $37.58\pm 0.03$ 77) in the literature. The DC relativistic contribution to $\alpha$ is $-0.018$ a.u., which is in excellent agreement with the value $-0.016$ a.u., as a correction added to the NR value, as shown in Table S26. There are no experimental values available for Be in the literature yet.
The SR and NR values (71.20 and 71.54) for magnesium (Mg, $Z=12$) are in excellent agreement with the CCSD(T) results (71.15 and 71.63) at the complete basis set limit in Ref. 78, where “0” in X2C-0/NR-0 means all electrons are correlated. This suggests that the basis set used in this work for Mg is sufficient to provide an accurate $\alpha$. With the same basis sets, the DC value (71.19) agrees well with the recommended value ($71.2\pm 0.4$) in Ref. 3. Additionally, the DC value also aligns with all experimental values and is close to all other computational values in Table S26.
The SR and NR values (157.61 and 159.90) of calcium (Ca, $Z=20$) are in excellent agreement with the CCSD(T) results (157.61 and 159.81) at the complete basis set limit reported in Ref. 78. In this context, “10” in X2C-10/NR-10, as shown in Table S26, indicates that 10 core electrons are treated using ECP. This suggests that the basis set used in this work for Ca is sufficient to provide an accurate $\alpha$. With the same basis sets, the DC value (157.55) agrees well with the recommended value ($160.8\pm 4$) in Ref. 3. Lim et al. computed DC and SR values for alkaline-earth-metal elements from Ca to radium (Ra, $Z=88$) using a similar procedure as in this work. However, the basis set used for Ca in their work is ($32s22p10d9f6g$), which is much larger than the largest basis set ($20s15p8d5f$) in Ref. 79. Additionally, larger basis sets are also employed for atoms from Sr to Ra in this work compared to Ref. 79. The SOC contribution to $\alpha$ is evaluated at the MBPT2 level in Ref. 79, while the corresponding values at CCSD and CCSD(T) levels are also available in this work, as shown in Table S2. This gives confidence that the DC value obtained in this work is more accurate than the value reported in Ref. 79.
The DC value ($197.67\pm 0.01$) is slightly higher than the upper bound of the recommended value ($197.2\pm 0.2$) from Ref. 3, corresponding to the computational values in Refs. 72, 80, where a hybrid relativistic CI and MBPT method is used. The difference could be attributed to the higher-order correlation contributions included in this work. In general, the DC/SR values (197.67/197.82) are close to all experimental values and computational values listed in Table S26. The NR value ($209.64\pm 0.01$) with the basis set ($34s26p16d8f6g2h$) also aligns with the result (210.64) with the basis set ($21s16p11d9f3g$) in Ref. 79, suggesting that the basis set used in this work is sufficient. A difference between the SR value (197.82) and the result (198.62) in Ref. 78 could be mainly attributed to the difference between all-electron and ECP methods.
It can be seen from Table 2 that the error bar for barium (Ba, $Z=56$) is still large based on the recommended value ($272.9\pm 10$) in Ref. 3. The difference between the DC value (273.51) and the CCSD(T) results ($273\pm 4.1$ 57, 79 and 272.8 81) lies within the uncertainty range suggested by Ref. 3. It should be noted that a larger basis set ($36s31p20d8f6g2h$) is employed compared to the one ($24s22p16d12f4g1h$) used in Ref. 81. The Gaunt contribution ($-0.01$ a.u.) is also studied by Borschevsky et al. at both the DHF and correlation stages. In addition, Dutta recently studied the trend of the Gaunt interaction contribution to $\alpha$ for alkaline-earth elements, finding the Gaunt contribution to be $-0.454$ a.u. at the level of 3rd-order perturbation theory.82 To be conservative, the latter value is used for uncertainty estimation.
The DC value of Ra (245.68) agrees well with the recommended value ($246\pm 4$) in Ref. 3 and the computational values ($246.2\pm 4.9$ 57, 79 and 242.8 81) obtained using a similar procedure as in this work. It should be noted that a larger basis set ($38s35p24d16f6g2h$) is employed compared to the one ($26s23p18d13f5g2h$) used in Ref. 81. The Gaunt contribution evaluated at the CCSD(T) level is tiny and negligible for Ra based on Ref. 81, while it is $-1.88$ a.u. when evaluated at the level of 3rd-order perturbation theory in Ref. 82. To be conservative, the latter value is used for uncertainty estimation.
4.2.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 2 elements. Considering the number of core electrons and virtual shells is similar as used for heavy atoms in the previous work,81 the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are not estimated explicitly. They are included implicitly in the error due to $\Delta P_{\text{basis}}$. Several SR test calculations were performed for basis sets using dyall.cv4z and s-aug-dyall.cv4z, compared to the results with d-aug-dyall.cv4z. The Gaunt contribution, based on Ref. 82, is considered the main source of the error due to $P_{\text{others}}$.
In SR-CC calculations of Be, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 36.93, 37.78, and 37.79 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. Therefore, the error due to the finite basis set, $\Delta P_{\text{basis}}$, can be approximated as the difference of SR values (0.1 a.u.) between this work and Ref. 78. Based on Ref. 82, the Gaunt contribution is less than 0.01 a.u. and negligible. Additionally, the error of 0.1 a.u. is also larger than the difference between the DC value and the recommended value in Ref. 3. In conclusion, the error of 0.1 a.u. in this study for Be is acceptable.
In SR-CC calculations of Mg, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 70.64, 71.20, and 71.20 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. Therefore, the error due to the finite basis set, $\Delta P_{\text{basis}}$, can be approximated as the difference of SR values (0.05 a.u.) between this work and Ref. 78. Based on Ref. 82, the Gaunt contribution is less than 0.02 a.u. Additionally, the difference between CCSD and CCSD(T) is $0.63$ a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation. In conclusion, the error of 0.32 a.u. in this study for Mg is acceptable.
In SR-CC calculations of Ca, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 156.19, 157.61, and 157.61 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. However, the difference in the most accurate SR values between this study and Ref. 78 is negligible. Therefore, the error due to the finite basis set, $\Delta P_{\text{basis}}$, is approximated as the difference in NR values (0.09 a.u.) between this work and Ref. 78. Based on Ref. 82, the Gaunt contribution is less than 0.03 a.u. Additionally, the difference between CCSD and CCSD(T) is $2.84$ a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation. In conclusion, the error of 1.42 a.u. in this study for Ca is acceptable.
In SR-CC calculations of Sr, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 197.45, 197.82, and 197.78 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. Considering the large difference between all-electron and ECP methods, the SR values in Ref. 78 are not used in uncertainty estimation for Sr, Ba, and Ra in this work. The SR CCSD(T) difference (0.04 a.u.) between s-aug-dyall.cv4z and d-aug-dyall.cv4z is approximately taken as the error due to the finite basis set. Based on Ref. 82, the Gaunt contribution is less than 0.04 a.u. Additionally, the difference between CCSD and CCSD(T) is $4.62$ a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation. In conclusion, the error of 2.31 a.u. in this study for Sr is acceptable.
In SR-CC calculations of Ba, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 272.83, 273.96, and 273.46 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. The SR CCSD(T) difference (0.50 a.u.) between s-aug-dyall.cv4z and d-aug-dyall.cv4z is approximately taken as the error due to the finite basis set. Based on Ref. 82, the Gaunt contribution is less than 0.46 a.u. Additionally, the difference between CCSD and CCSD(T) is $8.18$ a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation. In conclusion, the error of 4.15 a.u. in this study for Ba is acceptable.
In SR-CC calculations of Ra, as shown in Table S2, the results for CCSD(T) with dyall.cv4z, s-aug-dyall.cv4z, and d-aug-dyall.cv4z are 246.22, 247.69, and 247.20 a.u., respectively, suggesting that the convergence of $\alpha$ with the dyall.cv4z family basis set is reached. The SR CCSD(T) difference (0.49 a.u.) between s-aug-dyall.cv4z and d-aug-dyall.cv4z is approximately taken as the error due to the finite basis set. Based on Ref. 82, the Gaunt contribution is less than 1.88 a.u. Additionally, the difference between CCSD and CCSD(T) is $9.02$ a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation. In conclusion, the error of 4.91 a.u. in this study for Ra is acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 2.
4.2.3 The correlation and relativistic effects on polarizabilities
Figure 2(a) displays the relationship between dipole polarizabilities and atomic numbers for Group 2 elements. It reveals that non-relativistic polarizabilities increase monotonically with atomic numbers. However, the trend for relativistic polarizabilities resembles the non-relativistic case for lighter atoms but is lower for heavier elements, particularly for Sr and Ba, where deviations exceed 11 and 42, respectively. Moreover, a considerable decrease in relativistic polarizabilities is observed from Ba to Ra.
Figure 2(b) highlights that the scalar-relativistic effect is more significant than the SOC effect, leading to an approximate $Z^{2}$ increase in the relativistic effects on the scalar-relativistic contribution to dipole polarizabilities, paralleling the trend observed in Group 1 elements. The SOC contribution to dipole polarizabilities for Group 2 atoms is small.
Figure 2(c) indicates that the electron correlation contribution in all relativistic calculations grows monotonically in absolute value with increasing atomic numbers, except for Ra atom where the electron-correlation contributions of relativistic calculations are nearly 10 less than in the non-relativistic simulations.
Figure 2: Polarizabilities (in a.u.) of Group 2 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and full-relativistic Dirac-Coulomb (DC) dipole polarizabilities. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.3 Group 13 elements
4.3.1 Comparison with literature
The results computed by CCSD(T) and MRCISD are presented in Table 3, compared with the recommended values in Ref. 3. The component-resolved results for the ${{}^{2}}P_{3/2}$ state are also listed. Additional data can be found in Table S3. The results obtained by Eq. (15) can be found in Table S11, and the corresponding summary of the most accurate results is listed in Table S19.
Both DC-CC and DC-CI values in this work align well with the recommended theoretical values presented in Ref. 3, except for boron (B, $Z=5$) and nihonium (Nh, $Z=113$). It should be noted that the DC-CC data for B is not available in the present work due to convergence issues encountered with DHF using a DC Hamiltonian. Therefore, the most accurate results are DC-CC results for all Group 13 elements except for B, for which the SR value is used as the most accurate result. For Al, the best DC-CC result is obtained by including 11 outer-shell electrons and 223 virtual shells for electron correlation instead of all electrons and virtual shells, as shown in Table S3. Furthermore, all component-resolved values for Group 13 atoms are consistent with the results computed in Ref. 8. While an earlier study has shown that virtual cutoff values of 10 are sufficient for DC calculations,8 I opted for a tight cutoff of 25 in this work. This choice stems from the confidence in earlier findings and the anticipation that a higher cutoff might ensure augmented accuracy.
Table 3: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 13 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
B Al Ga In Tl Nh
$\hat{H}$ State Method
NR ${}^{2}P$ CCSD(T) $20.41$ $55.50\pm 0.47$ $52.76\pm 0.05$ $69.25\pm 0.05$ $70.61\pm 1.10$ $88.88\pm 0.70$
SRn ${}^{2}P$ CCSD(T) $20.40\pm 0.01$ $55.54\pm 0.43$ $52.30\pm 0.13$ $66.91\pm 0.20$ $64.37\pm 0.94$ $81.62\pm 0.11$
SRd ${}^{2}P$ CCSD(T) $20.40\pm 0.01$ $55.54\pm 0.43$ $52.00\pm 0.09$ $66.98\pm 0.07$ $64.53\pm 1.01$ $81.96\pm 0.11$
DC ${}^{2}P_{1/2}$ CCSD(T) --- - $57.51\pm 1.75$ $51.08\pm 0.32$ $61.92$ $51.26\pm 0.13$ $32.89\pm 0.18$
MRCISD $20.35$ $57.61\pm 0.02$ $50.66$ $62.87$ $51.26$ $28.21$
${}^{2}P_{3/2}$ MRCISD $20.38$ $58.07\pm 0.01$ $52.86$ $70.55$ $81.71$ $129.98\pm 0.01$
${}^{2}P_{3/2},M_{J}=1/2$ MRCISD $22.31$ $66.81\pm 0.01$ $63.64$ $84.75$ $106.41$ $188.76\pm 0.01$
${}^{2}P_{3/2},M_{J}=3/2$ MRCISD $18.46$ $49.33$ $42.08$ $56.36$ $57.01$ $71.21$
Rec. --- - --- - $20.40\pm 0.09$ $57.51\pm 2.47$ $51.08\pm 1.19$ $61.92\pm 1.92$ $51.26\pm 1.04$ $32.89\pm 1.60$
Ref. 3 --- - --- - $20.5\pm 1$ $57.8\pm 1.0$ $50.3\pm 3$ $65\pm 4$ $50.0\pm 2$ $29.2\pm 2$
The most accurate theoretical and experimental values for Group 13 elements are collected in Table S27. Next, a detailed comparison between this work and the results in the literature is presented, considering only the ${}^{2}P_{1/2}$ state. Only computational and experimental values close to the recommended values in Ref. 3 are compared. For comparison of scalar-relativistic polarizability, only the $\text{SR}_{n}$ value in this work is considered when $\text{SR}_{n}$ and $\text{SR}_{d}$ are different due to the fewer number of correlated electrons used in the latter calculations.
For B, the computed NR value is 20.41, closely matching literature values ($20.47$ 83, $20.43\pm 0.11$ 31, $20.42$ 75, $20.480$ 84, and $20.30\pm 0.26$ 85), all using the CCSD(T) method. The difference between the NR value in this work and the one in Ref. 84 could be attributed to the finite basis sets, and it can therefore be used for uncertainty estimation. The DC-CI value (20.35/20.38) for the ${}^{2}P_{1/2}/^{2}P_{3/2}$ state is slightly lower than the MRCI result (20.53/20.54 a.u.) reported by Fleig 8. This could be due to the triple excitations not being included and the larger basis set used in this work.
For Al, the DC-CC value ($57.51\pm 1.75$) is in good agreement with the recent CCSD result ($58.8\pm 1.2$) that includes Breit and QED contributions,86 and experimental results ($55.3\pm 5.5$ 87, 88, $54\pm 11$ 89, and $54.9\pm 5.3$ 90). The computed DC-CI value ($57.61\pm 0.02$) is also close to other CI results ($54.4\pm 2.2$ 8 and $57.8\pm 1.0$ 91). The calculated NR polarizability is $57.15\pm 0.08$, which is in good agreement with literature values such as $57.74$ by Lupinetti et al. 92 and $57.85$ by Manz et al. 75, using the CCSD(T) method. Additionally, the NR, SR, and DC values are $64.00\pm 0.89$, $64.12\pm 0.82$, and $56.74\pm 0.26$, respectively, obtained using Eq. (15), as shown in Table S19. The NR and SR values are much higher than computational values in the literature, suggesting that Eq. (15) is insufficient to obtain accurate $\alpha$ for Al.
For Ga, In, and Tl, the most accurate values in the literature for the ground state, to my knowledge, are reported by Borschevsky et al. using CCSD(T) methods.93 Therefore, I mainly focus on the comparison between this study and Ref. 93 for these three atoms. The DC-CC result ($51.08\pm 0.32$) for Ga is in excellent agreement with the result (51.05) reported by Borschevsky et al. using the CCSD(T) method.93 Additionally, the computed CCSD result (52.18), as shown in Table S3, is also close to the CCSD value (52.48) reported in Ref. 93. The negligible difference could be attributed to the different basis sets used, the number of electrons correlated, or the fitting procedure. The basis set ($31s22p14d6f4g1h$) with a total number of 322 functions is used in this work, while the basis set ($22s19p12d9f3g1h$) with a total number of 307 functions is used in Ref. 93. However, compared to Ref. 93 where 21 electrons are correlated, only 13 electrons are used in this work. Moreover, three data points are used for fitting $\alpha$ in Ref. 93, while five data points are employed in this study. The Gaunt contribution (0.18) is also evaluated in Ref. 93, which can be used for uncertainty estimation. The DC-CI result (50.66) agrees well with other CI results (49.9 8 and $51.3\pm 2.0$ 91). The computed NR polarizability for Ga is $52.76\pm 0.05$, which is consistent with values such as $52.91\pm 0.40$ by Cernusak et al. 94 and $53.01$ by Manz et al. 75. Compared to the experimental value, the DC-CC value is slightly higher than the upper bound of the experimental value ($46.6\pm 4.0$) reported by Ma et al.56
The DC-CC result (61.92) for In is in excellent agreement with the result (62.22) reported by Borschevsky et al. using the CCSD(T) method.93 Additionally, the computed CCSD result (60.17), as shown in Table S3, is also close to the CCSD value (62.31) reported in Ref. 93. The difference could be attributed to the different basis sets used, the number of electrons correlated, or the fitting procedure. The basis set ($34s28p19d6f4g1h$) with a total number of 373 functions is used in this work, while the basis set ($23s21p14d9f4g1h$) with a total number of 341 functions is used in Ref. 93. Compared to Ref. 93, where 21 electrons are correlated with four-component DC-CC, 31 electrons are correlated with two-component DC-CC in this work. Moreover, three data points are used for fitting $\alpha$ in Ref. 93, while five data points are employed in this study. The DC-CI result (62.87) agrees well with other CI results ($61.9\pm 1.2$ 8 and $66.4\pm 5.0$ 91). The DC-CC and DC-CI results are also in good agreement with the experimental value ($62.1\pm 6.1$).56
The DC-CC result (51.26) for Tl is slightly smaller than the result (52.14) reported by Borschevsky et al. using the CCSD(T) method.93 Additionally, the computed CCSD result (51.43), as shown in Table S3, is also slightly smaller than the CCSD value (52.56) reported in Ref. 93. It should be noted that 35 electrons are correlated with four-component DC-CC in both this work and Ref. 93. The difference can only be attributed to the different basis sets used or the fitting procedure. The basis set ($35s32p22d15f4g1h$) with a total number of 494 functions is used in this work, while the basis set ($26s24p18d13f5g2h$) with a total number of 453 functions is used in Ref. 93. Three data points are used for fitting $\alpha$ in Ref. 93, while five data points are employed in this study. The Gaunt contribution ($-0.72$) is also evaluated in Ref. 93, which can be used for uncertainty estimation. Another set of four-component DC-CC calculations were performed with only 13 electrons being correlated in this work. The corresponding CCSD(T) result is $51.94\pm 0.20$, suggesting that a difference of 0.68 a.u. is observed due to the smaller number of electrons correlated. Furthermore, the DC-CI result (51.26) agrees well with other CI results (51.6 8 and $50.7\pm 5.0$ 91). The DC-CC and DC-CI results are also in good agreement with the experimental value ($51.3\pm 5.4$).95
For Nh, the DC-CC value ($32.89\pm 0.18$) slightly exceeds the upper bounds of the recommended value ($29.2\pm 2$). Two earlier studies, FS-CCSD 9 and SD+CI,10 have reported values of 29.85 and 28.8, respectively. The CCSD result (32.38) is also higher than the recommended value (29.85) by Pershina et al. using FS-CCSD. The difference could be attributed to the different basis sets used, the number of electrons correlated, or the fitting procedure. The basis set ($36s36p25d17f4g1h$) with a total number of 545 functions is used in this work, while the basis set ($26s24p18d13f5g2h$) with a total number of 453 functions is used in Ref. 9. However, compared to Ref. 9 where 46 electrons are correlated, only 35 electrons are used in this work. Three data points are used for fitting $\alpha$ in Ref. 9, while five data points are employed in this study.
To further validate the computed data in this work, several tests are performed. First, a DC-CC calculation with 13 electrons being correlated is performed, and the CCSD(T) result is $33.11\pm 0.15$, as shown in Table S3. The result shows that the difference is only 0.22 a.u. Second, a two-component DC-CC calculation with 49 electrons being correlated is performed, and the CCSD(T) result is $32.79$, as shown in Table S3, suggesting that including more electrons only contributes 0.1 a.u. to $\alpha$. The last test is a two-component DC-CC calculation with a d-aug-dyall.acv4z basis set, i.e., ($38s38p27d19f6g3h$) with a total number of 657 functions, and 51 electrons correlated. Only the CCSD value is computed due to the extremely high computational cost. The CCSD result is $32.57\pm 0.02$, and only a 0.02 a.u. difference is observed compared to the CCSD value (32.59) in the second test. This implies that the basis sets and number of correlated electrons are converged in this work. It should be noted that the computed DC-CI value (28.21) for Nh in this study aligns well with the previous SD+CI result (28.8) presented in Ref. 10. This also implies that a new recommended value for Nh should be updated based on the CCSD(T) result in this work.
4.3.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 13 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are not estimated explicitly. They are included implicitly in the error due to $\Delta P_{\text{basis}}$.
For B, the difference between the computed NR value (20.41) and the result (20.480) in Ref. 84 is used as the error due to the finite basis set ($\Delta P_{\text{basis}}$). The differences ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in NR-CC and SR-CC calculations are 0.04 and 0.02 a.u., respectively, with the larger one taken as the error due to $\Delta P_{\text{(T)}}$. The contribution due to the full DC relativistic effect is about 0.05 a.u. between NR-CC and DC-CI for the ground state of B. This value is employed as the error due to $\Delta P_{\text{others}}$. In conclusion, the error of 0.09 a.u. in this study for B is acceptable.
For Al, the uncertainty of the DC value due to fitting Eq. (14) is 1.75 a.u., which is much higher than the uncertainty (1.0 a.u.) recommended in Ref. 3. Therefore, the error of 1.75 a.u. in this study for Al is acceptable.
For Ga, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is 1.1 a.u., which is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. The Gaunt contribution at the CCSD(T) level is 0.18 a.u. based on Ref. 93, which is therefore employed as the error due to $\Delta P_{\text{others}}$. The error due to the finite basis set is taken as 0.50% of the DC-CC value as in Ref. 93. The difference in DC CCSD values (0.03 a.u.) between this work and Ref. 93 is treated as the error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$. In conclusion, the error of 1.15 a.u. in this study for Ga is acceptable, which is close to the uncertainty of 1.5 a.u. reported in Ref. 93.
For In, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is 1.75 a.u., which is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. The Gaunt contribution at the CCSD(T) level is not available in Ref. 93, and 0.72 a.u. for Tl is employed as the error due to $\Delta P_{\text{others}}$ for In. The error due to the finite basis set is empirically taken as 0.50% of the DC-CC value as in Ref. 93. The difference in DC CCSD(T) values (0.08 a.u.) between this work and Ref. 93 is treated as the error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$. In conclusion, the error of 1.92 a.u. in this study for In is acceptable, which is close to the uncertainty of 1.9 a.u. reported in Ref. 93.
For Tl, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is 0.17 a.u., which is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. The Gaunt contribution at the CCSD(T) level is 0.72 a.u. based on Ref. 93, which is employed as the error due to $\Delta P_{\text{others}}$. The error due to the finite basis set is empirically taken as 0.50% of the DC-CC value as in Ref. 93. The difference in CCSD(T) values (0.68 a.u.) between DC-CC calculations correlating 13 and 35 electrons is treated as the error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$. In conclusion, the error of 1.03 a.u. in this study for Tl is acceptable, which is smaller than the uncertainty of 1.6 a.u. reported in Ref. 93.
For Nh, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is 0.51 a.u., which is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. The Gaunt contribution at the CCSD(T) level is not available in Ref. 93; thus, 1.50 a.u. is taken as the error due to $\Delta P_{\text{others}}$ for Nh. The difference in CCSD values (0.02 a.u.) between two-component DC-CC calculations correlating 49 and 51 electrons is treated as the error due to the finite basis set. The difference in CCSD(T) values (0.10 a.u.) between DC-CC calculations correlating 35 and 49 electrons is treated as the error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$. In conclusion, the error of 1.59 a.u. in this study for Nh is acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 3.
4.3.3 The correlation and relativistic effects on polarizabilities
Figure 3(a) displays the relationship between dipole polarizabilities and atomic numbers for Group 13 elements. Also, the $M_{L}$ component polarizabilities at the spin-free-relativistic level are presented. In general, polarizabilities increase with atomic numbers at the non-relativistic level, with the exception of Ga, which has a smaller screening of the nuclear charge by $d$ electrons, resulting in more compact valence shells and a reduced polarizability as reported in the previous study.8 From the component-resolved results in Figure 3(a), the decrease in polarizability is mainly caused by the $M_{L}=\pm 1$ state. This effect is also observed in the comparison between In and Tl, but it is smaller than in the Ga case at the non-relativistic level. The difference in $M_{L}=0$ states of In and Tl eliminates the increase in the corresponding $M_{L}=\pm 1$ components, resulting in a polarizability value for Tl that is close to that of In in non-relativistic calculations. However, at the scalar-relativistic level, this change cannot be attributed solely to $f$ electrons screening due to relativity. In the Dirac picture, the SOC effect significantly reduces the polarizabilities of Tl and Nh. As a result, the polarizability of Tl is close to that of Ga, and the result for Nh is lower than Al. This finding is consistent with previous work in Ref. 8.
Both relativistic effects and electron correlation have a reducing effect on the polarizabilities of Group 13 elements, as seen in Figures 3(b) and (c). The DC relativistic effects ($\Delta\alpha_{r}^{\text{DC}}$) roughly increase with atomic numbers to the power of 2, as well as through the SOC effect, which is the dominant factor compared to scalar-relativistic effects ($\Delta\alpha_{r}^{\text{SR}}$). An exception is Al, whose DC-CC calculation is not accurate. The smaller difference between $\bar{\alpha}_{J=3/2}$ and $\alpha_{J=1/2}$ for B and Al in Table 3 also suggests that SOC contribution can be disregarded entirely, which is consistent with previous research presented in Ref. 8.
In Figure 3(c), the electron correlation effects generally increase with increasing atomic numbers at the non-relativistic and scalar-relativistic levels, with the exception of Tl, whose $\Delta\alpha_{c}^{\text{NR}}$ and $\Delta\alpha_{c}^{\text{SR}}$ are slightly larger than those of In, respectively. In the Dirac picture, electron correlation plays a critical role for Ga and In, whereas its effects decrease with increasing atomic numbers starting with In due to the rising influence of the SOC effect.
Figure 3: Polarizabilities (in a.u.) of Group 13 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. The results of states $M_{L}=0$ and $M_{L}=\pm 1$ are indicated as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations, respectively. (b) Illustration of the influence of relativistic effects, including SR, SOC, and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.4 Group 14 elements
4.4.1 Comparison with literature
The results computed by CCSD(T) are presented in Table 4, compared with the recommended values in Ref. 3. Additional results are presented in Table S4. The results obtained by Eq. (15) can be found in Table S12, and the corresponding summary of the most accurate results is listed in Table S20.
Table 4: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 14 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
C Si Ge Sn Pb Fl
$\hat{H}$ State Method
NR ${}^{3}P$ CCSD(T) $11.66$ $37.07\pm 0.01$ $39.70\pm 0.16$ $55.48\pm 0.15$ $62.20\pm 0.09$ $77.92\pm 0.13$
SRn ${}^{3}P$ CCSD(T) $11.67$ $37.10\pm 0.01$ $39.53\pm 0.07$ $54.18\pm 0.13$ $58.49\pm 0.10$ $72.54\pm 0.15$
SRd ${}^{3}P$ CCSD(T) $11.67$ $37.10\pm 0.01$ $39.70\pm 0.03$ $54.30\pm 0.04$ $58.49\pm 0.10$ $72.54\pm 0.15$
DC ${}^{3}P_{0}$ CCSD(T) $11.73$ $37.00\pm 0.01$ $39.19\pm 0.02$ $51.96$ $46.53$ $30.77$
Rec. --- - --- - $11.73\pm 0.48$ $37.00\pm 0.63$ $39.19\pm 0.74$ $51.96\pm 1.01$ $46.53\pm 0.89$ $30.77\pm 0.90$
Ref. 3 --- - --- - $11.3\pm 0.2$ $37.3\pm 0.7$ $40.0\pm 1$ $53.0\pm 6$ $47.0\pm 3$ $31.0\pm 4$
All DC results are in good agreement with the recommended values in Ref. 3 except for carbon (C, $Z=6$). In addition, the most accurate theoretical and experimental values for Group 14 elements are collected in Table S28 to validate the results in this work. Next, a detailed comparison between this work and the results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. Additionally, Thierfelder et al. computed the static dipole polarizabilities for Group 14 elements using the same method as in this work.96 Their computational values are also in excellent agreement with the recommended values from Ref. 3. Therefore, a more detailed comparison between this work and Ref. 96 is presented. For comparison of scalar-relativistic polarizability, only the $\text{SR}_{n}$ value in this work is considered when $\text{SR}_{n}$ and $\text{SR}_{d}$ are different due to the fewer number of correlated electrons used in the latter calculations.
The DC value (11.73) of C is slightly larger than the upper bound of the result ($11.26\pm 0.20$) reported in Ref. 96. The difference could be attributed to the different basis sets and the number of data points computed. The basis set ($19s11p6d4f2g$) is used in this work compared to the ($13s7p4d3f2g$) used in Ref. 96. In addition, energy with an external electric field of $+0.0005$ a.u. is also added in this work compared to Ref. 96. Moreover, the difference between SR and DC values, as shown in Table 4, could be attributable to the small SOC effect for carbon. This suggests that $LS$ coupling may provide a better description than $jj$ coupling. The calculated SR polarizability of C (11.61) is close to the DK value (11.70) in Ref. 96 and the recent scalar-relativistic CCSD(T) result (11.524).97 Due to the small relativistic effects, numerous NR values ($11.67\pm 0.07$ 31, 11.63 75, $11.64\pm 0.15$ 85, 11.683 84, 11.324 84) computed by the CCSD(T) method have been reported in the literature. The computed NR value (11.66) in this work is in good agreement with these values, with the largest error being less than 2%.
The DC, SR, and NR CCSD(T) values of silicon (Si, $Z=14$) are $37.00\pm 0.01$, $37.10\pm 0.01$, and $37.07\pm 0.01$, respectively, and are very close to the corresponding results (37.28, 37.58, and 37.54) reported in Ref. 96. The small difference in DC values is mainly attributed to the larger basis set ($25s15p8d6f4g$) used in this work compared to the one ($17s12p4d3f2g$) used in Ref. 96. In addition, the NR value in this work is in excellent agreement with NR CCSD(T) results ($37.4\pm 0.1$ 98, $37.17\pm 0.21$ 92, $37.16$ 75, and $36.90\pm 0.63$ 85) reported in the literature, as shown in Table S28. The difference between the SR value and the recent scalar-relativistic CCSD(T) result (36.803) is less than 1.2%.
The DC, SR, and NR CCSD(T) values for germanium (Ge, $Z=32$) are $39.19\pm 0.02$, $39.53\pm 0.07$, and $39.70\pm 0.16$, respectively, and are very close to the corresponding results (39.33, 39.97, and 40.16) reported in Ref. 96. The small difference in DC values could be attributed to the different basis sets and the number of correlated electrons. Specifically, a larger basis set ($31s22p14d6f4g1h$) is used in this work compared to the one ($23s19p15d4f2g$) used in Ref. 96. However, only 22 electrons are correlated in this work compared to all 32 electrons in Ref. 96. In addition, the NR value in this work is in excellent agreement with the recent NR CCSD(T) result ($39.78$).75
The DC, SR, and NR CCSD(T) values for tin (Sn, $Z=50$) are $51.96$, $54.18\pm 0.13$, and $55.48\pm 0.15$, respectively, and are very close to the corresponding results (52.70, 55.31, and 56.34) reported in Ref. 96. The difference in DC values could be attributed to the different basis sets and the number of correlated electrons. Specifically, a larger basis set ($34s28p19d6f4g1h$) is used in this work compared to the one ($36s29p21d6f2g$) used in Ref. 96. Additionally, 32 electrons are correlated in this work compared to only 22 electrons in Ref. 96. It should be noted that the Dyall basis set is also used for Sn in Ref. 96. Therefore, the difference can be used to evaluate the error due to the finite basis set. Furthermore, the DC value in this work is also in good agreement with experimental values ($42.4\pm 11$ 96 and $67.5\pm 8.8$ 56).
The DC, SR, and NR CCSD(T) values for lead (Pb, $Z=82$) are $46.53$, $58.49\pm 0.10$, and $62.20\pm 0.09$, respectively, and are very close to the corresponding results (47.34, 60.25, and 63.90) reported in Ref. 96. The difference in DC values could be attributed to the different basis sets and the number of correlated electrons. Specifically, the basis set ($35s32p22d15f4g1h$) is used in this work compared to the one ($37s33p25d19f2g$) used in Ref. 96. Additionally, 46 electrons are correlated in this work compared to only 36 electrons in Ref. 96. The Dyall basis set is also used for Pb in Ref. 96. Therefore, the difference can be used to evaluate the error due to the finite basis set. Furthermore, the DC value in this work is also in good agreement with experimental values ($47.1\pm 7.1$ 96, 95 and $56\pm 18$ 56).
The DC, SR, and NR CCSD(T) values for flerovium (Fl, $Z=114$) are $30.77$, $72.54\pm 0.15$, and $77.92\pm 0.13$, respectively, and are very close to the corresponding results (31.49, 79.07, and 82.12) reported in Ref. 96. The difference in DC values could be attributed to the different basis sets and the number of correlated electrons. Specifically, a larger basis set ($36s36p25d17f4g1h$) is used in this work compared to the one ($32s31p24d18f3g$) used in Ref. 96. Additionally, 50 electrons are correlated in this work compared to only 36 electrons in Ref. 96.
4.4.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 14 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are not estimated explicitly; they are included implicitly in the error due to $\Delta P_{\text{basis}}$. Considering a very similar method is employed in Ref. 96, the main difference is due to the basis sets. Therefore, the difference in DC values is taken as the error due to $\Delta P_{\text{basis}}$ for all Group 14 elements. The $\Delta P_{\text{(T)}}$ is quite small at the DC CCSD(T) level for all Group 14 elements, and they are $0.10$, $-0.56$, $-0.72$, $-0.66$, $-0.02$, and $0.38$ for C, Si, Ge, Sn, Pb, and Fl, respectively. Therefore, the error due to $\Delta P_{\text{(T)}}$ and higher-order correlation is approximated as the absolute value of $\Delta P_{\text{(T)}}$. The Gaunt contribution is also reported in Ref. 96 at the DHF level, which is employed as the main contribution to the error due to $\Delta P_{\text{others}}$. In conclusion, the total errors are 0.48, 0.63, 0.74, 1.01, 0.89, and 0.90 for C, Si, Ge, Sn, Pb, and Fl, respectively. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 4.
4.4.3 The correlation and relativistic effects on polarizabilities
The relationship between dipole polarizabilities and atomic numbers for Group 14 elements is depicted in Figure 4(a). In general, all dipole polarizabilities increase with atomic numbers except for atoms from Sn to Fl due to the SOC effect. Figure 4(b) shows that $\Delta\alpha_{r}^{\text{SOC}}$ and $\Delta\alpha_{r}^{\text{DC}}$ increase approximately as $Z^{2}$ with atomic numbers, while $\Delta\alpha_{r}^{\text{SR}}$ increases linearly as $Z$ with respect to atomic numbers. Moreover, the main difference between $\alpha^{\text{NR}}$ and $\alpha^{\text{SR}}$ is from the $M_{L}=0$ component, and the difference increases with atomic numbers, especially Pb and Fl. For lighter atoms C, Si, and Ge, all results are very close, suggesting that relativistic effects are negligible.
As shown in Figure 4(c), electron correlation reduces the dipole polarizabilities except for the DC value of Fl, which agrees with the previous findings in Ref. 96. The large deviation between $\Delta\alpha_{c}^{\text{SR}}$ and $\Delta\alpha_{c}^{\text{DC}}$ could be attributed to the fact that the full-relativistic DC DHF wave function is not a proper reference wave function, leading to considerable correlation energy for Si, Ge, and Sn.
Figure 4: Polarizabilities (a.u.) of Group 14 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. The results of states $M_{L}=0$ and $M_{L}=\pm 1$ are indicated as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.5 Group 15 elements
4.5.1 Comparison with literature
The results computed by CCSD(T) and MRCISD are presented in Table 5, compared with the recommended values in Ref. 3. The component-resolved results for the ${{}^{4}}S_{3/2}$ state are also listed. Additional results are provided in Table S5. The results obtained by Eq. (15) can be found in Table S13, and the corresponding summary of the most accurate results is listed in Table S21.
Table 5: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 15 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
N P As Sb Bi Mc
$\hat{H}$ State Method
NR ${}^{4}S$ CCSD(T) $7.25$ $24.92$ $29.83\pm 0.01$ $43.78\pm 0.01$ $51.18\pm 0.01$ $63.63\pm 0.03$
SRn ${}^{4}S$ CCSD(T) $7.24$ $24.93$ $29.70\pm 0.01$ $42.96\pm 0.01$ $48.60\pm 0.01$ $60.65\pm 0.03$
SRd ${}^{4}S$ CCSD(T) $7.27$ $25.16$ $29.62\pm 0.01$ $43.05$ $48.78\pm 0.01$ $60.96\pm 0.03$
DC ${}^{4}S_{3/2}$ MRCISD $7.20$ $24.83$ $28.98$ $41.98$ $46.57$ $68.64$
${}^{4}S_{3/2},M_{J}=1/2$ MRCISD $7.20$ $24.83$ $29.00$ $42.22$ $51.42$ $97.20$
${}^{4}S_{3/2},M_{J}=3/2$ MRCISD $7.20$ $24.83$ $28.96$ $41.73$ $41.71$ $40.07$
Rec. --- - --- - $7.24\pm 0.15$ $24.93\pm 0.42$ $29.70\pm 0.66$ $41.98\pm 1.10$ $46.57\pm 2.23$ $68.64\pm 7.69$
Ref. 3 --- - --- - $7.4\pm 0.2$ $25.0\pm 1$ $30\pm 1$ $43\pm 2$ $48\pm 4$ $71.0\pm 20$
All DC, SR, and NR polarizabilities of the ground state for Group 15 elements are in excellent agreement with the recommended values reported in Ref. 3, except for the DC value of As (28.98), which is slightly smaller than the lower bound of the corresponding recommended value ($30\pm 1$). This implies that the relativistic effects are small for all Group 15 elements except for Mc, where a large uncertainty is used. Therefore, in this work, the most accurate calculations for nitrogen (N, $Z=7$), phosphorus (P, $Z=15$), and arsenic (As, $Z=33$) are SR CCSD(T), while DC MRCISD is used for antimony (Sb, $Z=51$), bismuth (Bi, $Z=83$), and Mc due to the increasing SOC contribution to polarizabilities.
In addition, the most accurate theoretical and experimental values for Group 15 elements are collected in Table S29 to validate the results in this work. Surprisingly, there are no DC values available for N, P, and As yet, based on Table S29. For Sb, only one DC value ($45\pm 11$) is reported by Doolen with relativistic local density approximation (LDA) in linear response theory 99 based on Table S29. Additionally, this method is also used to compute the DC value ($50\pm 12$) of Bi. The most accurate DC values for Bi and Mc in the literature are reported by Dzuba et al. using SD+CI methods. It should be noted that the DC values for Bi and Mc in this work are 46.57 and 68.64, respectively, which are in excellent agreement with the SD+CI values (44.62 and 70.5).10
Next, a detailed comparison between this work and the SR and NR results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. For computational data, only values computed with CCSD(T) methods in the literature are considered. For comparison of scalar-relativistic polarizability, only the $\text{SR}_{n}$ value in this work is considered when $\text{SR}_{n}$ and $\text{SR}_{d}$ are different due to the fewer number of correlated electrons used in the latter calculations.
The SR value (7.24) of N in this work is in excellent agreement with the experimental values ($7.6\pm 0.4$ 50, 100 and 7.28 101, 102). Additionally, the SR value is also close to another SR CCSD(T) result, 7.178, reported by Canal Neto et al.97 The NR value (7.25) is in good agreement with previous computational values, e.g., 7.21,75 $7.2$,103 and 7.25±0.09plus-or-minus7.250.097.25$\pm$ 0.097.25 ± 0.09.85 However, a slightly larger deviation ($<0.12$ a.u.) is observed between the NR value in this work and the values (7.367 and 7.153), including the extrapolation to the complete basis set limit, reported by Éhn and Černušák. This difference is therefore employed as the error due to the $\Delta P_{\text{basis}}$.
No experimental data is available for P in the literature for comparison. The SR value (24.93) in this work is in excellent agreement with the SR data (24.980) reported in Ref. 97. The NR value (24.92) also perfectly aligns with recent NR CCSD(T) results (24.88,75 25.0,103 $24.95\pm 0.43$ 85) where very large basis sets are used. This suggests that the Dyall basis set is also sufficient to obtain a reasonable dipole polarizability for P.
Similarly, no experimental data is available for As in the literature for comparison. The SR value ($29.70\pm 0.01$) in this work is in excellent agreement with the SR data (29.92) reported in Ref. 97, and the SR results obtained by the ECP method, e.g., 29.81 102 and 29.6.103 The NR value ($29.83\pm 0.01$) is also close to the counterpart (29.65) reported in Ref. 75.
Again, no experimental data is available for Sb in the literature for comparison. The SR value ($42.96\pm 0.01$) in this work is in excellent agreement with the SR data computed by CCSD(T) with the ECP method ($43.03$,102 $42.78$,75 and $42.8$ 103). Additionally, the NR value ($43.78\pm 0.01$) is slightly larger than the counterpart (42.26) reported in Ref. 104. This difference can be attributed to the much larger uncontracted basis set ($34s28p19d6f4g1h$) used in this work compared to the contracted one [$6s6p4d1f$] used in Ref. 104.
For Bi, the DC value lies within the uncertainty of the experimental value ($55\pm 11$).56 The SR value ($48.60\pm 0.01$) in this work is in excellent agreement with the SR data computed by CCSD(T) with the ECP method ($48.75$,102 $48.88$,75 and $48.8$ 103).
It should be noted that there are no NR and SR data available in the literature for Mc based on Table S29. In this work, they are computed to be 60.84 and $60.64\pm 0.02$, respectively, by CCSD(T).
4.5.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 15 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are approximated as the difference between $\text{SR}_{n}$ and $\text{SR}_{d}$. Based on the comparison with the literature, the Dyall basis set is sufficient. To be conservative, 0.5% of the most accurate results are used as the error due to $\Delta P_{\text{basis}}$, except for N, where 0.12 a.u. is taken.
Additionally, the $\Delta P_{\text{(T)}}$ in SR calculations is always positive and less than 0.1 a.u. for all elements except for Mc, where the $\Delta P_{\text{(T)}}$ is -0.39 a.u. Therefore, the difference between $\text{SR}_{d}$ and DC values is used as the error due to $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{others}}$. In conclusion, the total errors are 0.15, 0.42, 0.66, 1.10, 2.23, and 7.69 for N, P, As, Sb, Bi, and Mc, respectively. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 5.
4.5.3 The correlation and relativistic effects on polarizabilities
The relationship between dipole polarizabilities and atomic numbers for Group 15 elements is depicted in Figure 5(a). It can be seen that the dipole polarizability generally increases with atomic number. Additionally, the SOC effects are as crucial as scalar-relativistic effects starting from Sb, as shown in Figure 5(b). Both $\Delta\alpha_{c}^{\text{NR}}$ and $\Delta\alpha_{c}^{\text{SR}}$ increase with atomic numbers and reduce polarizabilities, starting from P. Additionally, relativistic effects generally reduce polarizabilities, with the exception of Mc, where the contribution of the SOC effect significantly improves the polarizability by nearly 8.0. Figure 5(c) suggests that the electron correlation effects are small in both non-relativistic and scalar-relativistic calculations.
Figure 5: Polarizabilities (a.u.) of Group 15 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.6 Group 16 elements
4.6.1 Comparison with literature
The results computed using CCSD(T) and MRCISD are presented in Table 6, compared with the recommended values in Ref. 3. The component-resolved results for the ${{}^{2}}P_{3/2}$ state are also listed. Additional results are provided in Table S6. The results obtained by Eq. (15) can be found in Table S14, and the corresponding summary of the most accurate results is listed in Table S22.
Table 6: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 16 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
O S Se Te Po Lv
$\hat{H}$ State Method
NR ${}^{3}P$ CCSD(T) $5.16$ $19.17\pm 0.03$ $25.62\pm 0.27$ $38.56\pm 0.28$ $45.83\pm 0.02$ $58.72\pm 0.01$
SRn ${}^{3}P$ CCSD(T) $5.15$ $19.21\pm 0.02$ $25.57\pm 0.19$ $38.01\pm 0.09$ $44.32$ $57.00\pm 0.01$
SRd ${}^{3}P$ CCSD(T) $5.16$ $19.35\pm 0.02$ $25.53\pm 0.15$ $37.94\pm 0.14$ $44.48$ $57.22\pm 0.01$
DC ${}^{3}P_{2}$ MRCISD $5.13$ $19.08$ $24.48$ $36.58$ $43.22$ $71.66$
${}^{3}P_{2},M_{J}=0$ MRCISD $5.41$ $20.21$ $25.87$ $38.00$ $43.83$ $68.41$
${}^{3}P_{2},M_{J}=1$ MRCISD $4.84$ $17.94$ $23.08$ $35.16$ $42.61$ $74.92$
${}^{3}P_{2},M_{J}=2$ MRCISD $5.27$ $19.64$ $25.18$ $37.29$ $43.52$ $70.04$
Rec. --- - --- - $5.15\pm 0.04$ $19.21\pm 0.32$ $25.57\pm 1.08$ $36.58\pm 1.37$ $43.22\pm 1.28$ $71.66\pm 14.45$
Ref. 3 --- - --- - $5.3\pm 0.2$ $19.4\pm 0.1$ $28.9\pm 1.0$ $38.0\pm 4$ $44.0\pm 4$-- -
The DC-CI results are in agreement with the recommended values presented in Ref. 3, except for sulfur (S, $Z=16$) and selenium (Se, $Z=34$). Considering the small relativistic effects for light elements, the most accurate calculations for oxygen (O, $Z=8$), S, and Se in this work are SR CCSD(T). In contrast, DC MRCISD is used for tellurium (Te, $Z=52$), polonium (Po, $Z=84$), and Lv due to the increasing SOC contribution to polarizabilities.
In addition, the most accurate theoretical and experimental values for Group 16 elements are collected in Table S30 to validate the results in this work. Next, a detailed comparison between this work and results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. For SR and NR results, only values computed with CCSD(T) methods in the literature are considered. In addition, for SR polarizability, only the $\text{SR}_{n}$ value in this work is considered when $\text{SR}_{n}$ and $\text{SR}_{d}$ are different due to the fewer number of correlated electrons used in the latter calculations.
The SR value (5.15) of O in this work is in excellent agreement with the experimental value ($5.2\pm 0.4$).100 The NR value (5.16) is in good agreement with previous computational values, e.g., $5.24\pm 0.04$,57, 31 $5.21\pm 0.07$,85 and $5.145$.84 The NR, SR, and DC values obtained by fitting Eq. (15) are 5.16, 5.16, and 5.13, respectively, suggesting that the difference due to the fitting procedure is small for O.
No experimental data is available for S in the literature for comparison. The NR value ($19.17\pm 0.03$) in Table 6 for S is slightly smaller than the recommended value ($19.4\pm 0.1$) and other computational values in Table S30, e.g., 19.37±0.12plus-or-minus19.370.1219.37$\pm$ 0.1219.37 ± 0.12.92 The NR, SR, and DC values, obtained by fitting Eq. (15), are $19.41\pm 0.02$, $19.44\pm 0.02$, and 19.10, respectively. This suggests that the precision of CCSD(T) energies should be improved to obtain valid NR and SR hyperpolarizabilities for S using Eq. (14).
The DC value for Se (24.48) is much lower than the recommended value ($28.90\pm 1.0$), which is based on the experimental result from Ref. 105. However, the scalar-relativistic value ($25.57\pm 0.19$) in this study closely aligns with previous computational values ($26.24\pm 0.52$ 106 and 26.65 107). The NR CCSD value (25.19) is close to the linear-response CCSD value (25.03).75 It should be noted that negative NR and SR hyperpolarizabilities are observed at the CCSD(T) level, leading to large uncertainties in $\alpha$ as shown in Table 6.
The SR CCSD value (37.63) for Te, as shown in Table S6, is close to the linear-response CCSD values (37.65 108 and 37.51 75). It should be noted that negative NR and SR hyperpolarizabilities are observed at the CCSD(T) level, leading to a large uncertainty in $\alpha$ for Te, as shown in Table 6. The difference between SR DHF and CCSD values is 0.35, as shown in Table S6, indicating a small contribution to $\alpha$ due to electron correlation for Te. Therefore, the NR and SR CCSD values (38.08 and 37.63) could be more accurate for Te than CCSD(T) due to large numerical errors in hyperpolarizability.
The SR value (44.32) of Po is close to the recent CCSD(T) value (44.77) reported by Centoducatte et al.109 The SR CCSD value (43.97) agrees well with the result (44.22) computed by CCSD with the ECP method.75
The theoretical values for Lv are also estimated, with the NR, SR, and DC values being $58.72\pm 0.01$, $57.00\pm 0.01$, and 71.66, respectively. Accounting for SOC effects in the Lv calculation led to a reduction of 14.66 in the dipole polarizability compared to the scalar-relativistic calculation. The scalar-relativistic results at the CCSD(T) level closely match its DHF values. This suggests that the influence of electron correlation is small when determining the dipole polarizability of Lv in scalar-relativistic calculations.
4.6.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 16 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are approximated as the difference between $\text{SR}_{n}$ and $\text{SR}_{d}$. Based on the comparison with the literature, the Dyall basis set is sufficient. To be conservative, 0.5% of the DC results in Table 6 are used as the error due to $\Delta P_{\text{basis}}$ for each atom. The difference between $\text{SR}_{d}$ and DC values is used as the error due to $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{others}}$. In conclusion, the total errors are 0.04, 0.32, 1.06, 1.37, 1.28, and 14.45 for O, S, Se, Te, Po, and Lv, respectively. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 6.
4.6.3 The correlation and relativistic effects on polarizabilities
The relationship between dipole polarizabilities and atomic numbers for Group 16 elements is presented in Figure 6(a). This figure also highlights the $M_{L}=0$ and $M_{L}=\pm 1$ polarizabilities, denoted as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations, respectively. It is observed that the polarizabilities slightly increase with atomic numbers, and a significant discrepancy between scalar-relativistic and DC polarizabilities occurs in Lv due to the SOC effect. From Figure 6(b), relativistic effects slightly reduce polarizabilities, but in the case of Lv, the SOC effects notably increase the polarizability. Lastly, Figure 6(c) shows that electron correlation marginally enhances polarizabilities in atoms up to Te but reduces them in heavier atoms beyond Te.
Figure 6: Polarizabilities (in a.u.) of Group 16 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. The results of states $M_{L}=0$ and $M_{L}=\pm 1$ are indicated as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations, respectively. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.7 Group 17 elements
4.7.1 Comparison with literature
The results computed using CCSD(T) and MRCISD are presented in Table 7, compared with the recommended values in Ref. 3. The component-resolved results for the ${{}^{2}}P_{3/2}$ and ${{}^{2}}P_{1/2}$ states are also listed. Additional results are provided in Table S7. The results obtained by Eq. (15) can be found in Table S15, and the corresponding summary of the most accurate results is listed in Table S23.
Table 7: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 17 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
F Cl Br I At Ts
$\hat{H}$ State Method
NR ${}^{2}P$ CCSD(T) $3.60$ $14.54$ $20.57\pm 0.22$ $33.20\pm 0.96$ $39.50$ $50.26\pm 0.12$
SRn ${}^{2}P$ CCSD(T) $3.61$ $14.56\pm 0.01$ $21.00\pm 0.30$ $32.06\pm 0.06$ $38.39$ $49.91\pm 0.35$
SRd ${}^{2}P$ CCSD(T) $3.60$ $14.68$ $20.69\pm 0.27$ $32.07\pm 0.06$ $38.48$ $50.06\pm 0.35$
DC ${}^{2}P_{3/2}$ MRCISD $3.57$ $14.36$ $20.06$ $31.28$ $38.92$ $65.15$
${}^{2}P_{3/2},M_{J}=3/2$ MRCISD $3.70$ $14.94$ $20.95$ $32.63$ $41.59$ $71.89$
${}^{2}P_{3/2},M_{J}=1/2$ MRCISD $3.43$ $13.78$ $19.16$ $29.92$ $36.25$ $58.40$
${}^{2}P_{1/2},M_{J}=1/2$ MRCISD $3.57$ $14.40$ $20.27$ $32.01$ $40.79$ $227.28$
Rec. --- - --- - $3.61\pm 0.04$ $14.56\pm 0.35$ $21.00\pm 0.77$ $31.28\pm 0.81$ $38.92\pm 0.49$ $65.15\pm 15.09$
Ref. 3 --- - --- - $3.74\pm 0.08$ $14.6\pm 0.1$ $21.0\pm 1$ $32.9\pm 1.3$ $42.0\pm 4$ $76.0\pm 15$
The DC values align well with the recommended values, except for the light atoms, F and chlorine (Cl, $Z=17$). Given the small relativistic effects in light elements, the most accurate calculations for F, Cl, and Br in this work are SR CCSD(T). In contrast, DC MRCISD is used for I, At, and Ts due to the increasing SOC contribution to polarizabilities. As expected, the polarizabilities of states $J=\frac{3}{2},M_{J}=\frac{1}{2}$ are found to be smaller than those of the state $J=\frac{3}{2},M_{J}=\frac{3}{2}$. Moreover, the values for $J=\frac{3}{2},M_{J}=\frac{3}{2}$ are also slightly larger than those for $J=\frac{1}{2},M_{J}=\frac{1}{2}$ across all elements except for Ts, which displays an anomalously high polarizability, nearly four times larger than that of the $\alpha_{3/2,1/2}$ state. This deviation can be attributed to the heavily relativistic contraction of the $\phi\left(\frac{1}{2},\frac{1}{2}\right)$ spinors, which are less polarizable and more occupied in the $J=\frac{3}{2}$ states as reported in Ref. 11. The small values of $\alpha_{J=1/2}-\bar{\alpha}_{J=3/2}$ and $\bar{\alpha}_{M_{L}}-\bar{\alpha}_{3/2}$ suggest that the SOC effect can be neglected in the calculation of dipole polarizabilities for Group 17 atoms, except for Ts, where the SOC effects are particularly significant.
In addition, the most accurate theoretical and experimental values for Group 17 elements are collected in Table S31 to validate the results in this work. It can be seen that there are no DC values available for Ts yet. Next, a detailed comparison between this work and results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. For SR and NR results, only values computed with CCSD(T) methods in the literature are considered. In addition, for SR polarizability, only the $\text{SR}_{n}$ value in this work is considered when $\text{SR}_{n}$ and $\text{SR}_{d}$ are different due to the fewer number of correlated electrons used in the latter calculations.
The NR value (3.60) of F in this work is slightly smaller than other computational results ($3.70\pm 0.03$ 31, $3.68\pm 0.05$ 85, $3.701$ 84, and $3.655$ 84) obtained by the CCSD(T) method. The NR value of 3.62 is obtained by using the same data points and numerical differentiation equation used in Ref. 84, suggesting that the deviation could be attributed to the finite basis set.
The NR value (14.54) of Cl is in excellent agreement with other theoretical results ($14.57\pm 0.10$ 92, $14.59\pm 0.25$ 85, and $14.582$ 97) computed with the CCSD(T) method.
The DC-CI value (20.06) of the ${}^{2}P_{3/2}$ state for Br is close to the result (21.7) obtained by Fleig et al. using the MRCI method.11 The NR CCSD value (20.45), as shown in Table S7, also aligns with the result (20.4) obtained by the linear-response CCSD method.75
The DC-CI value (31.28) of the ${}^{2}P_{3/2}$ state for I is slightly smaller than the lower bound of the recommended value ($32.9\pm 1.3$), which is based on the experimental value reported by Maroulis et al.57 Additionally, it is also slightly smaller compared to the MRCI value (34.1 11 and $33.0\pm 1.7$ 57, 11, 110). This deviation could be attributed to the lack of higher-order excitations or the larger basis set used in this work. The SR CCSD value (31.73), as shown in Table S7, is in good agreement with the result (31.57) obtained by the linear-response CCSD with the ECP method.
The DC-CI value (38.92) of the ${}^{2}P_{3/2}$ state for At is slightly smaller than the MRCI value (41.9) reported by Fleig et al.11 The SR CCSD(T) value (38.39) is close to CCSD(T) results (40.4 3 and 41.1 111) obtained by the ZORA scalar-relativistic CCSD(T). However, the SR CCSD value (38.06), as shown in Table S7, is in excellent agreement with the linear-response CCSD result (38.93) 108 and the CCSD value (38.15) computed with the ECP method.75
For Ts, only empirically estimated values are available in the literature.112, 12 This results in a very large uncertainty for Ts, as recommended in Ref. 3. In this work, NR, SR, and DC values are evaluated using systematic methods similar to those used for other elements in the same group. The final DC value is 65.15 for the ${}^{2}P_{3/2}$ state, which aligns well with the recommended value ($76.0\pm 15$).3
4.7.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 17 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are approximated as the difference between $\text{SR}_{n}$ and $\text{SR}_{d}$. Based on the comparison with the literature, the Dyall basis set is sufficient. To be conservative, 0.5% of the DC results in Table 7 are used as the error due to $\Delta P_{\text{basis}}$ for each atom. The difference between $\text{SR}_{d}$ and DC values is used as the error due to $\Delta P_{\text{(T)}}$ and $\Delta P_{\text{others}}$. In conclusion, the total errors are 0.04, 0.35, 0.71, 0.81, 0.49, and 15.09 for F, Cl, Br, I, At, and Ts, respectively. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 7.
4.7.3 The correlation and relativistic effects on polarizabilities
The relationship between dipole polarizabilities and atomic numbers for Group 17 elements is presented in Figure 7(a). This figure also highlights the $M_{L}=0$ and $M_{L}=\pm 1$ polarizabilities, denoted as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations, respectively. All results indicate that dipole polarizabilities increase with atomic numbers. Figure 7(b) provides the group trend of relativistic effects on the static polarizabilities of Group 17 atoms. It is observed that the SOC effect significantly improves polarizabilities for the element Ts, indicating that the SOC effects are particularly significant for Ts. The electron correlation contribution is presented in Figure 7(c), indicating that electron correlation is relatively small, with values less than 2 for non-relativistic calculations and less than 1 for scalar-relativistic calculations. For light atoms, electron correlation slightly increases polarizabilities at both non-relativistic and scalar-relativistic levels, whereas for heavy atoms, it reduces polarizabilities.
Figure 7: Polarizabilities (in a.u.) of Group 17 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. The results of states $M_{L}=0$ and $M_{L}=\pm 1$ are indicated as NR0 (SR0) and NR1 (SR1) for non-relativistic (scalar-relativistic) calculations, respectively. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
4.8 Group 18 elements
4.8.1 Comparison with literature
Table 8 summarizes the CCSD(T) results for Group 18 elements, compared with the recommended values from Ref. 3. Additional results can be found in Table S8. The results obtained by Eq. (15) can be found in Table S16, and the corresponding summary of the most accurate results is listed in Table S24.
Table 8: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 18 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
He Ne Ar Kr Xe Rn Og
$\hat{H}$ State Method
NR ${}^{1}S$ CCSD(T) $1.38$ $2.57$ $11.01$ $16.78$ $27.23\pm 0.01$ $34.08\pm 0.01$ $45.26\pm 0.02$
SR ${}^{1}S$ CCSD(T) $1.38$ $2.57$ $11.02$ $16.77$ $27.01$ $33.12$ $43.74$
DC ${}^{1}S_{0}$ CCSD(T) $1.38$ $2.57$ $11.02$ $16.80$ $27.20$ $35.07$ $58.24$
Rec. --- - --- - $1.38$ $2.57\pm 0.09$ $11.02\pm 0.06$ $16.80\pm 0.02$ $27.20\pm 0.14$ $35.07\pm 0.21$ $58.24\pm 0.32$
Ref. 3 --- - --- - $1.38$ $2.66$ $11.08$ $16.78\pm 0.02$ $27.32\pm 0.2$ $35.0\pm 2$ $58.0\pm 6$
The DC results for the heavy atoms, including krypton (Kr, $Z=36$), xenon (Xe, $Z=54$), radon (Rn, $Z=86$), and Og, show excellent agreement with the reference values in Ref. 3. For the lighter atoms, like Ne, and argon (Ar, $Z=18$), the polarizabilities calculated in this study are slightly lower than the recommended values as shown in Table 8. Moreover, the NR, SR, and DC values are 2.57 for Ne in this work, which is lower than experimental 113, 114, 2, 115 and theoretical values,116, 57, 117, 85, 118, 119 i.e., 2.66. Additionally, the SR and DC values are 11.02 for Ar in this work, which is lower than experimental 120, 121, 114, 57, 115 and theoretical values,57, 88, 92, 122 i.e., 11.08.
One possible reason for the discrepancy is the use of basis sets, which might not be sufficient to accurately describe the polarization in these light atoms. Another possible reason could be attributed to the use of the Gaussian model as the nucleus model, while a point nucleus model is usually employed in nonrelativistic programs.
In addition, the most accurate theoretical and experimental values for Group 18 elements are collected in Table S32 to validate the results in this work. Next, a detailed comparison between this work and results in the literature is presented. Only computational and experimental values close to the recommended values in Ref. 3 are compared. Only values computed with CCSD(T) methods in the literature are considered.
The NR, SR, and DC values are 1.38 for He in this work, which align well with the experimental values 120, 123, 1, 115 and other theoretical values.124, 125, 115, 126 The deviation is mainly attributed to the $\Delta_{\text{others}}$, including Breit and QED effects, among others.
The relativistic effects on dipole polarizabilities for Kr are small. Therefore, all NR, SR, and DC values (16.78, 16.77, and 16.80, respectively) for Kr are in good agreement with the experimental values ($16.766\pm 0.008$,121 $16.782\pm 0.005$,113 $16.740$,114, 127 and $16.734$ 114). The SR value also agrees well with the theoretical values ($16.80\pm 0.13$ 127 and 16.736 128) obtained using scalar-relativistic CCSD(T) methods in the literature. The DC value is also close to the recent MBPT value (16.800) including the Gaunt contribution.82
The DC value (27.20) for Xe is in good agreement with the experimental values (27.34227.34227.34227.342129 $27.078\pm 0.050$,113 and $27.292$ 114). The SR value (27.01) is also close to other computational values ($27.06\pm 0.27$,130 $27.2937\pm 0.0003$,116 $27.42$,131 $28.4\pm 0.5$,132 and $27.508$ 133).
There is no experimental data available for comparison for Rn in the literature. The DC value (35.07) for Rn is in excellent agreement with other computational DC results ($35.04\pm 1.8$,88, 9 $35.0$,134 $35.3$,135 and $35.53\pm 0.36$ 136) in the literature.
There is no experimental data available for comparison for Og in the literature. The DC value (58.24) for Og in this work is in excellent agreement with other computational DC results ($57.98$ 137 and $56.5\pm 1.8$ 136). Moreover, it is observed that the hyperpolarizability using Eq. (14) for Og is negative in this work. Therefore, the values fitted using Eq. (15) are also computed, and the NR, SR, and DC values are 45.28, 43.77, and $57.90\pm 0.06$, respectively, as shown in Table S24.
Jerabek et al. reported NR, SR, and DC values for Og, using the same method and basis set as in this work, as 45.30, 43.78, and 57.98, respectively. The Gaunt contribution is also included in the DC value. In Ref. 137, the numerical differentiation method is used, and only $\alpha$ is fitted.138 It can be seen that the NR and SR values fitted by Eq. (15) perfectly align with the counterparts reported in Ref. 137. The difference in DC values (0.08) could be mainly attributed to the Gaunt contribution.
Additionally, the 4C Hamiltonian is used in this work, while a 2C Hamiltonian is used in Ref. 137. To study the difference due to the 2C and 4C Hamiltonians used, three test calculations were performed. The first and second calculations used a 2C Hamiltonian with 18 and 40 electrons correlated, denoted as 2C-DC-CC@dyall.acv4z@(core 18)[vir 278] and 2C-DC-CC@dyall.acv4z@(core 40)[vir 278], respectively, as shown in Table S16. The CCSD(T) results fitted using Eq. (15) are 57.86 and 57.96, respectively. The DHF values for both cases are 60.58. The third test calculation used 2C-CC with 50 electrons correlated, and the Gaunt contribution to $\alpha$ was also evaluated at the DHF level, denoted as 2C-DC-Gaunt-CC@dyall.acv4z@(core 50)[vir 278] in Table S16. The DHF and CCSD(T) results fitted using Eq. (15) are 60.45 and 57.79, respectively. The difference between the CCSD(T) results of the third test and the CCSD(T) value reported in Ref. 137 is about 0.19 a.u., which could be attributed to the Gaunt contribution evaluated at the CCSD(T) level in Ref. 137. Comparing the results between the second and third calculations, the Gaunt contribution to $\alpha$ at the DHF level is about $-0.13$. Therefore, the total Gaunt contribution is about 0.06, which is close to the observed difference (0.08). These tests suggest that the difference between the 2C and 4C methods is negligible, and the Gaunt contribution to $\alpha$ is also small.
4.8.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 18 elements. For He, Ne, and Ar, accurate results are available in the literature. Therefore, the deviation between the DC value and the recommended value in Ref. 3 is taken as the uncertainty, which is mainly due to $\Delta P_{\text{others}}$.
Several SR test calculations with the d-aug-dyall.acv4z basis set are performed for Kr, Xe, Rn, and Og. Compared to the SR CCSD(T) values obtained with dyall.acv4z, the differences are about 0.01, 0.01, 0.01, and 0.13, respectively. The difference for Og with $\alpha$ obtained using Eq. (15) is 0.12. These values are used as the error due to the finite basis set ($\Delta P_{\text{basis}}$).
The difference between 2C and 4C Hamiltonians is studied for Ne, Ar, Xe, and Rn, as shown in Table S8, where the X2C methods with the same number of correlated electrons as in 4C-CC methods are performed. The differences are 0.00, 0.01, 0.01, and 0.03 a.u., respectively. The difference for Ar and Xe is 0.01; thus, it is assumed that it is also small for Kr in this work.
The Gaunt contribution has been investigated in Ref. 82, where Gaunt contributions to $\alpha$ are 0.0266, 0.0585, and 0.1813 for Kr, Xe, and Rn, respectively. It can be seen that the error due to 2C and 4C Hamiltonians is negligible. Therefore, only the Gaunt contribution is considered the main error due to $\Delta P_{\text{others}}$ for Kr, Xe, and Rn. For Og, the sum of the difference (0.33) due to using fitting equations defined in Eqs. (14) and (15) and the Gaunt contribution (0.08) is taken as the error due to $\Delta P_{\text{others}}$.
The error due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ is also studied for Kr, Xe, Rn, and Og. For Kr, a DC-CC calculation with 26 correlated electrons was performed compared to the most accurate calculation with 36 correlated electrons in this work; the differences at the CCSD and CCSD(T) levels are both 0.02.
For Xe, a 2C-DC calculation with 8 correlated electrons was performed compared to a 2C-DC calculation with 50 correlated electrons; the differences at the CCSD and CCSD(T) levels are 0.43 and 0.40, respectively. It should be noted that this difference is close to the difference (0.35) between the DHF and CCSD(T) values in the 2C-DC calculation with 50 correlated electrons. This implies that 50 correlated electrons are sufficient for $\alpha$ calculations for Xe.
Similarly, for Rn, a 2C-DC calculation with 18 correlated electrons was performed compared to another 2C-DC calculation with 50 correlated electrons. The difference is about 0.01, suggesting that 50 correlated electrons are sufficient for Rn.
For Og, a difference of 0.43 is obtained when two 2C-DC CC calculations with 18 and 40 correlated electrons were performed. However, this difference is reduced to 0.01 when comparing two DC-CC calculations with 40 and 50 correlated electrons.
Moreover, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in the DC-CC calculation is taken as the error due to $\Delta P_{\text{(T)}}$ and truncating higher-order correlation. They are 0.14, 0.20, 0.26, and 0.06 for Kr, Xe, Rn, and Og, respectively. The value for Og with the $\alpha$ obtained by Eq. (15) is 0.04. In conclusion, the errors of 0.14, 0.21, 0.32, and 0.43 for Kr, Xe, Rn, and Og, respectively, are acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 8.
4.8.3 The correlation and relativistic effects on polarizabilities
The relationship between dipole polarizabilities and atomic numbers for Group 18 elements is depicted in Figure 8(a), indicating that both non-relativistic and relativistic values increase with atomic numbers. The contributions of relativistic effects are depicted in Figure 8(b). The relativistic effects are small in light atoms like He, Ne, Ar, and Kr. As expected, the contribution of the SOC effect increases with atomic numbers, starting from Xe (0.20), increasing at Rn (1.95), and reaching a maximum for the Og atom (14.50). This finding suggests that the contribution of the SOC effect is indispensable for polarizability calculations of heavy atoms, especially for Og. Figure 8(c) displays the electron correlation effects, which are generally small ($<0.5$) for Group 18 elements. However, Og is an exception with $\Delta\alpha_{c}^{\text{DC}}=-2.36$.
Figure 8: Polarizabilities (in a.u.) of Group 18 elements. (a) Comparison of non-relativistic (NR), scalar-relativistic (SR), and Dirac-Coulomb (DC) dipole polarizabilities. (b) Illustration of the influence of relativistic effects, including SR, spin-orbit coupling (SOC), and DC, on dipole polarizabilities. (c) Examination of the impact of electron correlation on dipole polarizabilities in the presence of various relativistic effects.
5 Summary
In this study, I obtained the static dipole polarizabilities of main-group elements using the finite-field method combined with relativistic CCSD(T) and MRCI calculations. I also provide a systematic analysis of the impact of various relativistic effects on atomic dipole polarizabilities, including scalar-relativistic effects, spin-orbit coupling (SOC), and the full Dirac-Coulomb relativistic effect. For heavy atoms ($Z>40$), my theoretical values align well with the recommended values. Moreover, the analysis suggests that scalar-relativistic effects are predominant for elements in Groups 1–2, whereas SOC effects are typically negligible in these calculations. For elements in Groups 13–18, scalar-relativistic effects contribute minimally to dipole polarizability. However, the influence of SOC becomes significant starting from the fourth row of the Periodic Table for elements in Groups 13–14 and from the fifth row for elements in Groups 15–18. Additionally, the influence of electron correlation in the presence of various relativistic effects is investigated. Generally, it is found to be crucial for elements in Groups 1–2 and significant for those in Groups 13–14. However, for elements in Groups 15–18, its contribution is small. This study, employing computationally accurate methods with extensively large basis sets, yields comprehensive and consistent dipole polarizabilities for main-group elements, particularly for heavy atoms where only empirical values exist.
{acknowledgement}
Y.C. acknowledges the Foundation of Scientific Research - Flanders (FWO, file number G0A9717N) and the Research Board of Ghent University (BOF) for their financial support. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government. The author thanks Prof. Dr. Toon Verstraelen, and Dr. Thiago Carvalho Corso for helpful comments on the manuscript.
{suppinfo}
The Supplementary Material includes a PDF document and a file in “CSV” format. The PDF document contains the following information:
  1. 1.
    Additional detailed atomic dipole polarizabilities obtained by fitting Eq. (14), including those for DHF, MP2, and CCSD, as well as component-resolved results for open-shell atoms.
  2. 2.
    The atomic dipole polarizabilities obtained by fitting Eq. (15) are also included. The summary tables for the most accurate results for each group are provided.
  3. 3.
    The reference atomic dipole polarizabilities from literature, reported in Refs. 3, 13, are sorted by publication year for each group.
The CSV document contains the specified external electric field used for each $\alpha$ listed in the PDF document.